### v2

# The Adaptivity Barrier in Batched Nonparametric Bandits: Sharp Characterization of the Price of Unknown Margin

Abstract
We study batched nonparametric contextual bandits under a margin condition when the margin parameter $\alpha$ is unknown. To capture the statistical cost of this ignorance, we introduce the regret inflation criterion, defined as the ratio between the regret of an adaptive algorithm and that of an oracle knowing $\alpha$. We show that the optimal regret inflation grows polynomially with the horizon $T$, with exponent $\psi_{M}^{\star}$ given by the value of a convex optimization problem that depends on the dimension, smoothness, and number of batches $M$. Moreover, the minimizer of this optimization problem directly prescribes the batch allocation and exploration strategy of a rate-optimal algorithm. Building on this principle, we develop RoBIN (RObust batched algorithm with adaptive BINning), which achieves the optimal regret inflation up to polylogarithmic factors. These results reveal a new adaptivity barrier: under batching, adaptation to an unknown margin parameter inevitably incurs a polynomial penalty, sharply characterized by a variational problem. Remarkably, this barrier vanishes once the number of batches exceeds order $\log\log T$; with only a doubly logarithmic number of updates, one can recover the oracle regret rate up to polylogarithmic factors.
1 Introduction
A central question in sequential decision making is the cost of adaptation: how much performance is lost when key complexity parameters are unknown. Nonparametric contextual bandits provide a canonical setting to study this question (Woodroofe, 1979; Yang and Zhu, 2002; Rigollet and Zeevi, 2010; Perchet and Rigollet, 2013). In the fully online regime, the problem is well understood. Under smoothness and margin assumptions, algorithms that attain minimax-optimal regret can even adapt to an unknown margin parameter at no extra cost. In particular, the foundational work of Rigollet and Zeevi (2010) and the ABSE policy of Perchet and Rigollet (2013) demonstrate that margin adaptation comes at no statistical cost in the online regime.
In many settings, including clinical trials, education, and digital platforms, fully online interaction is infeasible because of logistical, ethical, or computational constraints. Instead, data collection proceeds in a limited number of batches: actions are fixed for a group of covariates, feedback is revealed only at the end of the batch, and subsequent policies must adapt accordingly. While minimax-optimal rates have been established for batched nonparametric contextual bandits when the margin parameter is known, these procedures require oracle knowledge to tune batch sizes and exploration schedules (Jiang and Ma, 2025). This raises a fundamental question: What is the statistical price of not knowing the margin parameter when learning under batch constraints?
This paper provides a sharp answer. We introduce the regret inflation criterion, defined as the ratio between the regret of an adaptive algorithm and that of an oracle who knows the true margin parameter. We show that the optimal regret inflation grows polynomially as $T^{\psi_{M}^{\star}}$ with the horizon $T$, with an exponent $\psi_{M}^{\star}$ precisely characterized by a convex variational optimization problem that depends on the dimension, smoothness, and batch budget. Strikingly, the minimizer of this program also prescribes the batch allocation and exploration schedule of a rate-optimal algorithm, yielding matching upper and lower bounds up to polylogarithmic factors.
A key corollary of this result is the identification of an adaptivity barrier unique to the batch constraint. In the online regime, the margin parameter admits free adaptation, but when updates are limited, adaptation becomes inherently costly. We prove that the barrier vanishes when the number of batches exceeds order $\log\log T$; that is, with a doubly logarithmic number of updates, one can match the oracle regret rate up to polylogarithmic factors. Conversely, when the batch budget grows more slowly than $\log\log T$, the regret inflation is unavoidably polynomial. This threshold cleanly delineates the transition between the regimes where batching constrains adaptation and where it does not.
Conceptually, our analysis unifies statistical limits and algorithm design through a single variational object. The variational characterization exposes the precise dependence of adaptivity cost on dimension, smoothness, and batch budget, while its minimizer yields an explicit constructive principle for designing robust batched policies under unknown complexity.
1.1 Related work
Contextual bandits.
The concept of contextual bandits was introduced by Woodroofe (1979). For linear contextual bandits, Auer (2002); Abbasi-Yadkori et al. (2011); Goldenshluger and Zeevi (2013); Bastani and Bayati (2020); Qian et al. (2023) established regret guarantees in both low- and high-dimensional settings. Meanwhile, modeling the mean reward function as a smooth function of the contexts was studied in Yang and Zhu (2002). Rigollet and Zeevi (2010) proved a minimax regret lower bound for this setup and designed an upper-confidence-bound-type (UCB-type) policy to attain the near-optimal rate. Perchet and Rigollet (2013) refined this result by proposing the Adaptively Binned Successive Elimination (ABSE) policy that can also adapt to the unknown margin parameter in the fully online setting. Additional insights in nonparametric contextual bandits were obtained in Qian and Yang (2016a); Reeve et al. (2018); Guan and Jiang (2018); Hu et al. (2022); Suk and Kpotufe (2021); Gur et al. (2022); Cai et al. (2024).
Margin condition in classification.
The margin condition originates from nonparametric classification, where it was studied by Audibert and Tsybakov (2007). This condition, often referred to as the Tsybakov margin condition, quantifies how well-separated the optimal decision boundary is and directly governs learning rates. Its adaptation to contextual bandits was initiated by Goldenshluger and Zeevi (2009); Rigollet and Zeevi (2010); Perchet and Rigollet (2013), who showed that the margin parameter $\alpha$ fundamentally shapes the complexity of bandit learning. In the online setting, adaptation to unknown $\alpha$ is possible without additional regret cost (Perchet and Rigollet, 2013). However, under the batch constraint, as explored in our work, such adaptivity becomes costly, giving rise to a new barrier.
Batch learning.
The multi-armed bandit problem under the batched setting was studied by Perchet et al. (2016); Gao et al. (2019). Batch learning in linear contextual bandits was studied by Han et al. (2020); Ren et al. (2020); Ruan et al. (2021) and Ren and Zhou (2023); Wang and Cheng (2020); Fan et al. (2023) further considered the problem with high-dimensional covariates. Jiang and Ma (2025); Arya (2025) studied the nonparametric contextual bandit problem under the batch constraint. Karbasi et al. (2021); Kalkanli and Ozgur (2021) developed batched Thompson sampling algorithms. Feng et al. (2022) considered the Lipschitz continuum-armed bandit problem under the batched setting. Further insights in batched bandits were developed in Zhang et al. (2020a); Jin et al. (2021a, b); Liu et al. (2025). Another related topic is online learning with switching costs (Cesa-Bianchi et al., 2013). Best arm identification with limited rounds of interaction has been studied by Tao et al. (2019). Reinforcement learning with low switching costs has been considered by Bai et al. (2019); Zhang et al. (2020b); Gao et al. (2021); Wang et al. (2021); Qiao et al. (2022).
Adaptation.
In the fully online setting, adaptivity to the margin parameter is feasible at no extra cost (Perchet and Rigollet, 2013). One might ask whether the same is true for the smoothness parameter. The answer is negative: even without the batch constraint, adaptation to smoothness is impossible (Locatelli and Carpentier, 2018; Gur et al., 2022; Cai and Pu, 2022). Minimax regret rates depend explicitly on the Hölder smoothness $\beta$, and no single procedure can achieve the optimal rate simultaneously across different values of $\beta$. This impossibility parallels classical results in nonparametric estimation and classification, where smoothness adaptation requires additional structure or necessarily incurs a penalty (Qian and Yang, 2016b; Gur et al., 2022; Cai et al., 2024). Thus, the margin parameter is the quantity of genuine interest: it admits free adaptation online, yet, becomes costly under the batch constraint.
1.2 Notation and paper organization
[…]
In terms of paper organization, in Section 2, we introduce the problem setup and the regret inflation criterion. Section 3 presents the main results, including the variational characterization of regret inflation and the master theorem. Section 4 describes the optimal algorithm guided by the variational principle. Section 5 contains the proof of the lower bound. We conclude with a discussion of extensions and future directions in Section 6.
2 Problem setup and the regret inflation criterion
We begin by introducing the model and assumptions for batched contextual bandits, then review the oracle regret when the margin parameter is known, and finally introduce the regret inflation criterion that drives the rest of our analysis.
2.1 Model and assumptions
We study a two-armed nonparametric contextual bandit with horizon $T$. At each round,
$(X_{t},Y_{t}^{(1)},Y_{t}^{(-1)}),\qquad t=1,\dots,T,$
are drawn i.i.d. from a distribution $P$, where the context $X_{t}\in\mathcal{X}\coloneqq[0,1]^{d}$ follows a distribution $P_{X}$. The rewards take values in $[0,1]$ and satisfy
$\mathbb{E}\!\big[Y_{t}^{(k)}\,|\,X_{t}\big]=f^{(k)}(X_{t}),\qquad k\in\{1,-1\},$
for unknown mean reward functions $f^{(1)},f^{(-1)}$.
Batch policies.
Under an $M$-batch constraint, the learner specifies (i) a partition $\Gamma=\{0=t_{0}<t_{1}<\cdots<t_{M}=T\}$ of the horizon, and (ii) a sequence of decision rules $\pi=(\pi_{t})_{t=1}^{T}$. At time $t$, only contexts up to $t$ and rewards from previous batches are available. Let $\Gamma(t)$ denote the batch index of round $t$, i.e., $\Gamma(t)\coloneqq i$ if $t_{i-1}<t\leq t_{i}$. The information set at time $t$ is
$\mathcal{H}_{t}=\{X_{\ell}\}_{\ell=1}^{t}\cup\{Y_{\ell}^{(\pi_{\ell}(X_{\ell}))}\}_{\ell=1}^{\,t_{\Gamma(t)-1}}.$
The grid $\Gamma$ may be chosen adaptively, meaning that the statistician can use all information up to $t_{i-1}$ to determine $t_{i}$. The statistician’s policy $\pi_{t}$ at time $t$ may depend on $\mathcal{H}_{t}$. The goal is to design an $M$-batch policy $(\Gamma,\pi)$ that can compete with an oracle that knows the environment, i.e., the law $P$ of $(X_{t},Y_{t}^{(1)},Y_{t}^{(-1)})$.
Distributional assumptions.
We impose the following standard conditions in the nonparametric bandits literature (Rigollet and Zeevi, 2010; Perchet and Rigollet, 2013; Cai et al., 2024; Jiang and Ma, 2025):
The first assumption is concerned with the regularity of the covariate distribution $P_{X}$.
Assumption 1 (Bounded density).
There exist constants $\underline{c},\bar{c}>0$ such that
$\underline{c}\,r^{d}\;\leq\;P_{X}\!\big(B(x,r)\big)\;\leq\;\bar{c}\,r^{d},\qquad\forall x\in\mathrm{supp}(P_{X}),\forall r\in(0,1],$
where $B(x,r)$ is the $\ell_{\infty}$ ball centered at $x$ with radius $r$.
The second assumption is on the smoothness of the mean reward functions.
Assumption 2 (Smoothness).
Each $f^{(k)}$ is $(\beta,L)$-Hölder smooth:
$|f^{(k)}(x)-f^{(k)}(x^{\prime})|\leq L\|x-x^{\prime}\|_{2}^{\beta},\qquad\forall x,x^{\prime}\in\mathcal{X},\;k\in\{1,-1\}.$
The last assumption measures the closeness between the reward functions of the two actions.
Assumption 3 (Margin).
For some $\alpha\geq 0$, there exist $\delta_{0}\in(0,1)$ and $D_{0}>0$ such that
$P_{X}\!\left(0<\big|f^{(1)}(X)-f^{(-1)}(X)\big|\leq\delta\right)\leq D_{0}\,\delta^{\alpha},\qquad\forall\delta\in[0,\delta_{0}].$
For a fixed margin parameter $\alpha$, let $\mathcal{P}_{\alpha}$ be the class of distributions satisfying Assumptions 13, where we implicitly assume that $d$ and $\beta$ are fixed and known.
Assumption 3 pertains to the margin condition in nonparametric classification (Mammen and Tsybakov, 1999; Tsybakov, 2004; Audibert and Tsybakov, 2007), and has been adapted to the bandit setup by Goldenshluger and Zeevi (2009); Rigollet and Zeevi (2010); Perchet and Rigollet (2013). The margin parameter governs the fundamental complexity of the problem. When $\alpha=0$, the margin assumption becomes vacuous, and the reward functions of the two arms can be arbitrarily close to each other, making it challenging to identify the optimal one. When $\alpha$ increases, the reward functions of the two actions exhibit strong separation over a region of high probability mass, and discerning the optimal action is less difficult.
The following proposition adapted from Perchet and Rigollet (2013) depicts the interplay between the smoothness parameter $\beta$ and the margin parameter $\alpha$.
Proposition 1.
Under Assumptions 1-3:
  • When $\alpha>d/\beta$, there is a constant gap between the reward functions of the two arms and one can take $\alpha=\infty$.
  • When $\alpha\leq d/\beta$, there exist nontrivial contextual bandit instances in $\mathcal{P}_{\alpha}$.
In other words, $\alpha>d/\beta$ is the regime where the problem class is reduced to multi-armed bandits without covariates and one equivalently has $\alpha=\infty$. On the other hand, $\alpha\leq d/\beta$ is the regime where $\mathcal{P}_{\alpha}$ corresponds to a non-degenerate class of nonparametric bandits.
2.2 Oracle regret with known margin
Given an $M$-batch policy $(\Gamma,\pi)$ and an environment $P$ with reward functions $(f^{(1)},f^{(-1)})$, we define the cumulative regret
$R_{T}(\Gamma,\pi;P)\coloneqq\mathbb{E}_{P}\!\left[\sum_{t=1}^{T}\Big(f^{\star}(X_{t})-f^{(\pi_{t}(X_{t}))}(X_{t})\Big)\right],$ (1)
where $f^{\star}(x)\coloneqq\max_{k\in\{1,-1\}}f^{(k)}(x)$ is the maximum mean reward one could obtain on the context $x$.
The oracle minimax regret with known $\alpha$ is
$R_{T}^{\star}(\alpha)\coloneqq\inf_{(\Gamma,\pi)}\;\sup_{P\in\mathcal{P}_{\alpha}}\;R_{T}(\Gamma,\pi;P),$ (2)
where $\mathcal{P}_{\alpha}$ denotes the class of environments with margin $\alpha$. It is known that the rate depends on the dimension $d$, smoothness $\beta$, margin $\alpha$, and batch budget $M$. Define
$\gamma(\alpha)\coloneqq\frac{\beta(\alpha+1)}{2\beta+d},\qquad\text{and}\qquad h_{M}(\alpha)\coloneqq\frac{1-\gamma(\alpha)}{1-\gamma(\alpha)^{M}}.$ (3)
When $\alpha=\infty$, we have $\gamma(\infty)=\infty$, and $h_{M}(\infty)=0$. The optimal rates with known margin have been established in Jiang and Ma (2025), which are stated in the following proposition.
Proposition 2.
Fix a margin $\alpha\in[0,d/\beta]\cup\{\infty\}$. We have
$c_{1}M^{-4}T^{\,h_{M}(\alpha)}\leq R_{T}^{\star}(\alpha)\leq c_{2}M(\log T)T^{\,h_{M}(\alpha)},$ (4)
where $c_{1},c_{2}>0$ are constants independent of $T$ and $M$.
Although Jiang and Ma (2025) established the minimax rate for the case $\alpha\leq 1/\beta$, we extend their result to $\alpha\in[0,d/\beta]\cup\{\infty\}$ for $d\geq 1$. Importantly, we make the dependence on $M$ explicit in the lower bound. See Appendix A for the proof.
2.3 The regret inflation criterion
When the margin parameter is unknown, the key question is: how much additional regret must we pay to adapt? To capture this, we introduce the notion of regret inflation.
Definition 1.
Denote by $\mathcal{K}\coloneqq[0,d/\beta]\cup\{\infty\}$ the set of possible margin parameters. For any $M$-batch policy $(\Gamma,\pi)$, define the regret inflation as
$\mathrm{RI}(\Gamma,\pi)\coloneqq\sup_{\alpha\in\mathcal{K}}\sup_{P\in\mathcal{P}_{\alpha}}\frac{R_{T}(\Gamma,\pi;P)}{R_{T}^{\star}(\alpha)}.$ (5)
This ratio compares the regret of the adaptive policy to that of the oracle who knows $\alpha$. Our notion of regret inflation parallels the risk inflation criterion in model selection for linear models (Foster and George, 1994). In that context, risk inflation measures the excess risk of a data-driven model selection rule relative to an oracle that knows the true model. Analogously, regret inflation quantifies the excess regret of an adaptive bandit policy relative to an oracle that knows the margin parameter. Both notions capture the statistical price of adaptation to unknown complexity.
Our goal is to characterize the rate of the optimal regret inflation,
$\inf_{\Gamma,\pi}\mathrm{RI}(\Gamma,\pi),$
and its dependence on the batch budget $M$.
3 Sharp characterization of the optimal regret inflation
The central task of this section is to analyze the optimal regret inflation. We show that it admits an exact characterization: its exponent is given by the value of a convex optimization problem, and the minimizers of this problem prescribe the design of the rate-optimal algorithm. Thus, the variational problem provides both a fundamental statistical limit and a constructive principle for algorithm design.
3.1 Key variational problem
Minimizing regret inflation can be formulated as a two-player zero-sum game. The learner commits to an $M$-batch policy $(\Gamma,\pi)$ without knowledge of the margin parameter, while nature selects a distribution $P$ over contexts and rewards consistent with some margin parameter $\alpha$. The payoff is the ratio between the learner’s regret and that of the oracle who knows $\alpha$.
Although both the learner’s strategy space (policies) and nature’s strategy space (distributions) are infinite-dimensional, the complexity of this game can be captured by a finite-dimensional reduction. The learner’s choice reduces to specifying a batch allocation across the $M$ updates, and nature’s choice reduces to selecting a margin parameter $\alpha$ from the admissible range $\mathcal{K}$. We parameterize the batch schedule $\Gamma=\{t_{0}=0<t_{1}<\cdots<t_{M}=T\}$ by exponents $u_{i}\in[0,1]$, so that $t_{i}\approx T^{u_{i}}$. This exponent parameterization captures exponentially growing batch sizes, which balance exploration and exploitation under geometric time scales. We refer to $\bm{u}=(u_{1},\dots,u_{M-1})$ as an exponent grid. This reduction yields a finite-dimensional convex optimization problem.
Formally, let $\bm{u}\in\mathcal{U}_{M}\coloneqq\{\bm{u}\in\mathbb{R}^{M-1}:0\leq u_{1}\leq\cdots\leq u_{M-1}\leq 1\}$ be the grid choice of the learner, and let $\alpha\in\mathcal{K}$ be the margin parameter selected by nature. Define the payoff function
$\Psi_{M}(\bm{u},\alpha)\;\coloneqq\;\max_{1\leq i\leq M}\eta_{i}(\bm{u},\alpha)-h_{M}(\alpha),$ (6)
where
$\eta_{1}(\bm{u},\alpha)=u_{1},\quad\eta_{i}(\bm{u},\alpha)=u_{i}-u_{i-1}\gamma(\alpha),\quad 2\leq i\leq M-1,\quad\eta_{M}(\bm{u},\alpha)=1-u_{M-1}\gamma(\alpha).$ (7)
Then the optimal value of this finite-dimensional two-player game is
$\psi_{M}^{\star}\coloneqq\inf_{\bm{u}\in\mathcal{U}_{M}}\sup_{\alpha\in\mathcal{K}}\;\Psi_{M}(\bm{u},\alpha).$ (8)
For notational convenience, we also define the objective function w.r.t. $\bm{u}$ as
$\psi_{M}(\bm{u})\coloneqq\sup_{\alpha\in\mathcal{K}}\Psi_{M}(\bm{u},\alpha).$ (9)
3.2 Optimal regret inflation
Our main theorem establishes a tight characterization of the optimal regret inflation.
Theorem 1.
Let $\psi_{M}^{\star}$ be the optimal value of the variational problem as defined in (8). Then there exist constants $c_{1},c_{2}>0$ independent of $T$ and $M$ such that
  1. 1.
    For any $M$-batch policy $(\Gamma,\pi)$,
    $\mathrm{RI}(\Gamma,\pi)\;\geq\;c_{1}M^{-8}\,(\log T)^{-1}\,T^{\psi_{M}^{\star}}.$
  2. 2.
[…]
    $\mathrm{RI}(\hat{\Gamma},\hat{\pi})\;\leq\;c_{2}\,M^{5}(\log T)\,T^{\psi_{M}^{\star}}.$
To better interpret Theorem 1, we treat $M$ as fixed and focus on how the regret inflation scales with the horizon $T$;the effect of varying $M$ will be analyzed in Section 3.4. Theorem 1 shows that the variational value $\psi_{M}^{\star}$ fully determines this dependence: up to logarithmic factors, any adaptive policy incurs inflation of order $T^{\psi_{M}^{\star}}$. If $\psi_{M}^{\star}$ is strictly positive, then Theorem 1 implies an unavoidable polynomial adaptation cost in $T$. The next subsection analyzes the structure of the variational problem and establishes that $\psi_{M}^{\star}$ is indeed strictly positive.
3.3 Structural properties of the variational problem
Proposition 3.
The following properties hold for the variational problem:
  1. 1.
    The function $\psi_{M}$ is convex and admits a minimizer $\bm{u}^{\star}$ in the interior of $\mathcal{U}_{M}$ with positive optimal value, i.e., $\psi_{M}^{\star}>0$.
  2. 2.
    There exists a non-increasing sequence $\{\alpha_{i}\}_{1\leq i\leq M}$ such that
    $\psi_{M}(\bm{u}^{\star})=\;\eta_{i}\bigl(\bm{u}^{\star},\alpha_{i}\bigr)\;-h_{M}(\alpha_{i}).$
    Moreover, one has $\alpha_{1}=\infty$, and $\alpha_{i}\leq d/\beta$ for $2\leq i\leq M$.
Part 1 of Proposition 3 ensures the exponent is strictly positive, implying polynomial inflation (up to polylogs) in the fixed $M$ case via Theorem 1. In Part 2 of Proposition 3, the sequence $\{\alpha_{i}\}$ identifies the branches of $\Psi_{M}$ that are active at $\bm{u}^{\star}$. This structure is used in the lower-bound construction (Section 5) to select difficulty levels of the hard instances. See Appendix B for the proof.
3.4 How many batches suffice for no regret inflation?
When we allow the number $M$ of batches to increase with the horizon $T$, a natural question arises: how many batches are sufficient for no regret inflation?
The following proposition provides a crude order-wise bound on $\psi_{M}^{\star}$ in terms of the number $M$ of batches.
Proposition 4.
Let $\gamma_{\mathrm{max}}\coloneqq\frac{d+\beta}{d+2\beta}$. Then we have the following control on the optimal exponent
$\frac{\gamma_{\mathrm{max}}^{M-1}(1-\gamma_{\mathrm{max}})^{2}}{(1-\gamma_{\mathrm{max}}^{M})^{2}}\leq\psi_{M}^{\star}\leq\frac{(M+1)\gamma_{\mathrm{max}}^{M-1}}{(1-\gamma_{\mathrm{max}})^{2}}.$
See Appendix B for the proof.
As a direct consequence of the upper bound in Proposition 4, we know that when $M\geq c_{1}\log\log T$ for some large constant $c_{1}$, we have $T^{\psi_{M}^{\star}}=O(1)$. Therefore, we know that $\log\log T$ batches are sufficient for no regret inflation. In other words, even if we do not know the exact margin parameter $\alpha$ of the bandit instance, with $\log\log T$ batches, there exists an algorithm that enjoys the same order of regret as the oracle algorithm with the knowledge of the $\alpha$, up to the $(\log\log T)^{5}\log T$ factor. This is summarized in the following corollary.
Corollary 1.
There exists a constant $c_{1}$ such that if $M\geq c_{1}\log\log T$, then
$\inf_{\Gamma,\pi}\;\mathrm{RI}(\Gamma,\pi)\leq C(\log\log T)^{5}\log T,$
where $C$ is some constant independent of $T$.
Conversely, by the lower bound in Proposition 4, we also know that $M\asymp\log\log T$ is necessary for small regret inflation. That is, when $M\leq c_{2}\log\log T$ for some small constant $c_{2}$, the regret inflation is super-polylogarithmic in $T$. Together, $M\asymp\log\log T$ marks a sharp phase transition between regimes with costly and cost-free margin adaptation.
3.5 Numerical illustrations
Figure 1: $\psi_{M}^{\star}$ vs. batch budget $M$ when $\beta=1$.
Figure 2: $\psi_{M}^{\star}$ vs. batch budget $M$ when $d=1$.
Since $\psi_{M}^{\star}$ generally lacks a closed-form expression, we first turn to numerical solutions. These experiments shed light on how the difficulty of adaptation varies with smoothness, dimension, and batch budget.
Figure 2 fixes $\beta=1$ and varies $d$. It plots the trend of $\psi_{M}^{\star}$ as the number $M$ of batches increases. Similarly, Figure 2 fixes $d=1$, and varies $\beta$. It also plots the trend of $\psi_{M}^{\star}$ as the number $M$ of batches increases. A few observations are in order.
  1. 1.
    All the exponents $\psi_{M}^{\star}$ are strictly positive, indicating a polynomial regret inflation when the margin parameter is not known.
  2. 2.
    When fixing $\beta$ and $M$, the exponent $\psi_{M}^{\star}$ is increasing in $d$. This demonstrates that adapting to $\alpha$ is increasingly difficult for high-dimensional problems.
  3. 3.
    Similarly, when fixing $d$ and $M$, the exponent $\psi_{M}^{\star}$ is increasing as $\beta$ decreases. This demonstrates that adapting to $\alpha$ is increasingly difficult for non-smooth problems.
In fact, the last two observations can be understood from the definition of the variational problem. Recall that under the setup for both plots, we have $\mathcal{K}=[0,d/\beta]\cup\{\infty\}$. Correspondingly, the allowable range for $\gamma(\alpha)$ is
$\gamma\in\left[\frac{\beta}{2\beta+d},\frac{d+\beta}{2\beta+d}\right]\cup\{\infty\}.$
It is clear that as $d$ increases and $\beta$ decreases, the allowable range for $\gamma$ increases. Since the inner maximization problem is effectively over $\gamma$, the supremum is non-decreasing. As a result, the optimal exponent $\psi_{M}^{\star}$ is non-decreasing.
4 The RoBIN algorithm: optimal adaptation to unknown margin
We now introduce RoBIN (RObust batched algorithm with adaptive BINning) to tackle batched contextual bandits with unknown margin. The algorithm builds on the BaSEDB framework from Jiang and Ma (2025), but introduces a new design principle: the batch schedule and split factors are selected based on the solution to the key variational convex program (8). As we shall see, this design principle enables robust adaptation across all margin parameter values.
We first describe the BaSEDB framework with fixed grid size and split factors, then present our new robust choices informed by the variational problem, and finally state the regret inflation guarantee for RoBIN.
Algorithm 1 RObust batched algorithm with adaptive BINning (RoBIN)
0: Batch size $M$, grid $\Gamma=\{t_{i}\}_{i=0}^{M}$, split factors $\{g_{i}\}_{i=0}^{M-1}$ as in Equations (11) and (12)
[…]
2:for each $C\in\mathcal{L}$ do
3:  $\mathcal{I}_{C}=\mathcal{I}$.
4:end for
5:for $i=1$ to M1M-1 do
6:  for $t=t_{i-1}+1$ to $t_{i}$ do
7:   $C\leftarrow\mathcal{L}(X_{t})$.
8:   Pull an arm from $\mathcal{I}_{C}$ in a round-robin way.
9:  end for
10:  Observe the outcomes for batch $i$
11:  Update $\mathcal{L}$ and $\{\mathcal{I}_{C}\}_{C\in\mathcal{L}}$ by invoking Algorithm 2 with inputs $(\mathcal{L},\{\mathcal{I}_{C}\}_{C\in\mathcal{L}},i,g_{i})$.
12:end for
13:for $t=t_{M-1}+1$ to $T$ do
14:  $C\leftarrow\mathcal{L}(X_{t})$.
15:  Pull any arm from $\mathcal{I}_{C}$.
16:end for
4.1 The BaSEDB framework with fixed parameters
We begin by reviewing the BaSEDB algorithm, which serves as the foundation for our robust variant. The BaSEDB algorithm operates over a horizon $T$ divided into $M$ batches, indexed by $i=1,\ldots,M$. It consists of three main components:
  • A batch schedule, specified by a grid $\Gamma=\{t_{0}=0<t_{1}<\cdots<t_{M}=T\}$, which determines the number of time steps in each batch;
  • A sequence of split factors $\{g_{i}\}_{i=0}^{M-1}$, which control how the covariate space $0,1]^{d}$ is iteratively partitioned into bins;
  • A Successive Elimination (SE) subroutine (see Algorithm 2) that runs independently in each active bin to eliminate suboptimal arms.
The high-level idea behind BaSEDB.
On a high level, initially, the covariate space is divided into $g_{0}^{d}$ bins of equal width. In each batch, the algorithm pulls arms uniformly at random within each bin and uses SE to discard clearly suboptimal arms. After the $i$-th batch, bins with unresolved ambiguity are split into finer bins using the corresponding split factor $g_{i}$, yielding a refined partition. The final batch is reserved for exploitation, where the algorithm chooses arms based on the surviving set in each bin; see Algorithm 1.
A detailed description of BaSEDB.
More formally the BaSEDB algorithm can be described using a tree diagram. Let $\mathcal{T}$ be a tree of depth $M$. The $i$-th level of the tree $\mathcal{T}$ consists of a collection of bins that is a regular partition of the covariate space $\mathcal{X}$, which we denote by $\mathcal{B}_{i}$. Each bin $C\in\mathcal{B}_{i}$ has the same width $w_{i}$ given by $w_{0}=1$ and
$w_{i}\coloneqq(\prod_{l=0}^{i-1}g_{l})^{-1},\quad i\geq 1.$ (10)
Here $\{g_{i}\}_{i=0}^{M-1}$ is a list of split factors. More precisely, $\mathcal{B}_{i}$ is composed of all the bins
$C_{i,\bm{v}}=\{x\in\mathcal{X}:(v_{j}-1)w_{i}\leq x_{j}<v_{j}w_{i},1\leq j\leq d\},$
where $\bm{v}=(v_{1},v_{2},\ldots,v_{d})\in[\frac{1}{w_{i}}]^{d}$. Clearly, $\mathcal{B}_{i}$ has $(1/w_{i})^{d}$ bins in total.
Algorithm 1 operates in batches and keeps track of the following key variables: a collection $\mathcal{L}$ of active bins, and the set of active arms $\mathcal{I}_{C}$ for each $C\in\mathcal{L}$. The collection of active bins $\mathcal{L}$ is initialized to be $\mathcal{B}_{1}$, while the set of active arms $\mathcal{I}_{C}$ is set to be $\{1,-1\}$ for all $C\in\mathcal{L}$ at the beginning. During the $i$-th batch, each arm in $\mathcal{I}_{C}$ is pulled for an equal amount of times. At the end of that batch, the active arm set $\mathcal{I}_{C}$ is updated by doing a hypothesis testing based on the revealed rewards from this batch. If after the arm elimination process $|\mathcal{I}_{C}|>1$ for some $C\in\mathcal{L}$, this means the current bin $C$ is too coarse to distinguish the optimal action. Consequently, this bin $C$ is further split into its children $\textrm{child}(C,g_{i})$ in tree $\mathcal{T}$, which is a set of $g_{i}^{d}$ bins, and the child nodes $\textrm{child}(C,g_{i})$ will replace the original bin $C$ in $\mathcal{L}$. For the last batch, we simply pull any arm from $\mathcal{I}_{C}$ whenever the covariate $X_{t}$ lands in some $C\in\mathcal{L}$. [^1] For the final batch $M$, the split factor $g_{M-1}=1$ by default because there is no need to further partition the nodes for estimation.
Algorithm 2 Node splitting and arm elimination procedure
0: Active bin list $\mathcal{L}$, active arm sets $\{\mathcal{I}_{C}\}_{C\in\mathcal{L}}$, batch number $i$, split factor $g_{i}$.
[…]
2:for each $C\in\mathcal{L}$ do
3:  if $|\mathcal{I}_{C}|=1$ then
4:   $\mathcal{L}^{\prime}\leftarrow\mathcal{L}^{\prime}\cup\{C\}$.
5:   Proceed to next $C$ in the iteration.
6:  end if
7:  $\bar{Y}_{C,i}^{\max}\leftarrow\max_{k\in\mathcal{I}_{C}}\bar{Y}_{C,i}^{(k)}$.
8:  for each $k\in\mathcal{I}_{C}$ do
9:   if $\bar{Y}_{C,i}^{\max}-\bar{Y}_{C,i}^{(k)}>U(m_{C,i},T,C)$ then
10:    $\mathcal{I}_{C}\leftarrow\mathcal{I}_{C}-\{k\}$
11:   end if
12:  end for
13:  if $|\mathcal{I}_{C}|>1$ then
14:   for each $C^{\prime}\in\textrm{child}(C,g_{i})$ do
15:    $\mathcal{I}_{C^{\prime}}\leftarrow\mathcal{I}_{C}$.
16:   end for
17:   $\mathcal{L^{\prime}}\leftarrow\mathcal{L^{\prime}}\cup\textrm{child}(C,g_{i})$.
18:  else
19:   $\mathcal{L}^{\prime}\leftarrow\mathcal{L}^{\prime}\cup\{C\}$.
20:  end if
21:end for
22: Return $\mathcal{L}^{\prime}$.
Next, we turn to the arm elimination part in Algorithm 2. The underlying idea is based on Successive Elimination (SE) from the bandit literature (Even-Dar et al., 2006; Perchet and Rigollet, 2013; Gao et al., 2019). An arm is eliminated from the active arm set $\mathcal{I}_{C}$ if the revealed rewards from this batch provide sufficient evidence of the suboptimality of this arm. For any node $C\in\mathcal{T}$, denote by $m_{C,i}\coloneqq\sum_{t=t_{i-1}+1}^{t_{i}}\mathbf{1}\{X_{t}\in C\}$ the number of times the covariates go into the bin $C$ during the $i$-th batch. For $k\in\{1,-1\}$, define the empirical estimate arm $k$’s reward in bin $C$ during batch $i$ as
$\bar{Y}_{C,i}^{(k)}\coloneqq\frac{\sum_{t=t_{i-1}+1}^{t_{i}}Y_{t}\cdot\mathbf{1}\{X_{t}\in C,A_{t}=k\}}{\sum_{t=t_{i-1}+1}^{t_{i}}\mathbf{1}\{X_{t}\in C,A_{t}=k\}}.$
The expectation of $\bar{Y}_{C,i}^{(k)}$ is equal to
$\bar{f}_{C}^{(k)}\coloneqq\mathbb{E}[f^{(k)}(X)\mid X\in C]=\frac{1}{P_{X}(C)}\int_{C}f^{(k)}(x)\mathrm{d}P_{X}(x).$
A key quantity for SE is the uncertainty level of the estimates in bin $C$, which is given by
$U(\tau,T,C)\coloneqq 4\sqrt{\frac{\log(2T|C|^{d})}{\tau}},$
where $|C|$ is the width of the bin. The uncertainty level is defined in a way so that with high probability the suboptimal arm for bin $C$ is eliminated while the near optimal ones remain in it. If $|\mathcal{I}_{C}|>1$, then the surviving arms are statistically close to each other; we further split $C$ into finer bins to obtain a more precise estimate of the rewards of these actions in later batches.
4.2 Robust parameter design via variational optimization
In the original BaSEDB algorithm, both the batch schedule $\Gamma$ and split factors $\{g_{i}\}$ are chosen assuming knowledge of the true margin parameter $\alpha$, which allows the algorithm to achieve minimax-optimal regret in that setting.
To achieve optimal regret without knowledge of $\alpha$, RoBIN selects the split factors and the batch points based on the minimizer $\bm{u}^{\star}$ to the convex program (8).
Split factor design.
Based on $\bm{u}^{\star}$, we define the split factors as
$g_{0}=\lfloor T^{\frac{1}{2\beta+d}\cdot u_{1}^{\star}}\rfloor,\qquad\text{and}\qquad g_{i}=\lfloor T^{\frac{1}{2\beta+d}(u_{i+1}^{*}-u_{i}^{\star})}\rfloor,\qquad\qquad i=1,\dots,M-2.$ (11)
Batch grid construction.
In addition, we choose the grid to satisfy
$t_{i}-t_{i-1}$ $=\lfloor l_{i}w_{i}^{-(2\beta+d)}\log(Tw_{i}^{d})\rfloor,\quad 1\leq i\leq M-1,$ (12)
for some $l_{i}>0$ sufficiently large. Here, we recall that $w_{i}$ is given in Equation (10). It can be shown that under these choices, we have
$t_{i}\approx T^{u_{i}^{\star}},\quad 1\leq i\leq M-1,$
where the approximation sign ignores log factors.
To summarize, RoBIN runs the BaSEDB procedure using these robust parameters $(\Gamma,\{g_{i}\})$ specified in Equations (11) and (12).
4.3 Regret inflation guarantee
We now state the main guarantee for RoBIN.
Theorem 2.
[…]
$\mathrm{RI}(\hat{\Gamma},\hat{\pi})\apprle M^{5}(\log T)\cdot T^{\psi_{M}^{\star}}.$
This upper bound on the amount of regret inflation matches the lower bound result in Theorem 1 (up to $\log$ factors). It demonstrates that RoBIN could achieve the optimal adaptation cost when the margin parameter is unknown.
As we will soon see in its proof, the key to achieve optimal regret inflation is the use of the minimizer $\bm{u}^{\star}$ of the variational problem.
4.4 Proof of Theorem 2
We now prove that RoBIN achieves the optimal regret inflation rate. The argument proceeds in two steps. First, we establish a regret bound for the BaSEDB algorithm with an arbitrary grid $\bm{u}$ in the interior of $\mathcal{U}_{M}$. Second, we specialize to the minimizer $\bm{u}^{\star}$ of the variational problem (8), which yields the optimal rate.
Step 1: Regret of BaSEDB with a fixed grid.
Let $\bm{u}$ be any interior point of $\mathcal{U}_{M}$, i.e., $0<u_{1}<u_{2}<\cdots<u_{M-1}<1$. Consider the split factors and the grid size given by Equations (11) and (12). The following lemma provides a generic regret bound.
Lemma 1.
Fix any $\alpha\geq 0$, and let $\hat{\pi}_{\bm{u}}$ be as above. Then
$\sup_{P\in\mathcal{P}_{\alpha}}R_{T}(\hat{\Gamma}_{\bm{u}},\hat{\pi}_{\bm{u}};P)$ $\leq c\left(t_{1}+\sum_{i=2}^{M-1}(t_{i}-t_{i-1})\cdot w_{i-1}^{\beta+\alpha\beta}+(T-t_{M-1})w_{M-1}^{\beta+\alpha\beta}\right),$
where $c$ depends only on $(\beta,d)$.
See Appendix D for the proof.
Applying the relations (11)-(12), this bound simplifies to
$\sup_{P\in\mathcal{P}_{\alpha}}R_{T}(\hat{\Gamma}_{\bm{u}},\hat{\pi}_{\bm{u}};P)\;\lesssim\;(\log T)\Big(T^{u_{1}}+\sum_{i=2}^{M-1}T^{u_{i}-u_{i-1}\gamma(\alpha)}+T^{1-u_{M-1}\gamma(\alpha)}\Big).$
Since the maximum exponent dominates, we obtain
$\sup_{P\in\mathcal{P}_{\alpha}}R_{T}(\hat{\Gamma}_{\bm{u}},\hat{\pi}_{\bm{u}};P)\;\lesssim\;(\log T)M\,T^{\max_{1\leq i\leq M}\eta_{i}(\bm{u},\alpha)}.$
Dividing by the oracle regret $R_{T}^{\star}(\alpha)\asymp T^{h_{M}(\alpha)}$ from Proposition 2 yields
$\sup_{P\in\mathcal{P}_{\alpha}}\frac{R_{T}(\hat{\Gamma}_{\bm{u}},\hat{\pi}_{\bm{u}};P)}{R_{T}^{\star}(\alpha)}\;\lesssim\;M^{5}(\log T)T^{\max_{1\leq i\leq M}\eta_{i}(\bm{u},\alpha)-h_{M}(\alpha)}.$ (13)
Taking the supremum of (13) over all $\alpha\in\mathcal{K}$ gives
$\sup_{\alpha\in\mathcal{A}}\sup_{P\in\mathcal{P}_{\alpha}}\frac{R_{T}(\hat{\Gamma}_{\bm{u}},\hat{\pi}_{\bm{u}};P)}{R_{T}^{\star}(\alpha)}\;\lesssim\;M^{5}(\log T)\,T^{\sup_{\alpha\in\mathcal{A}}\big(\max_{1\leq i\leq M}\eta_{i}(\bm{u},\alpha)-h_{M}(\alpha)\big)}.$
Step 2: optimizing over $\bm{u}$.
Finally, by construction, RoBIN corresponds to the policy $\hat{\pi}_{\bm{u}^{\star}}$ where $\bm{u}^{\star}$ minimizes the right-hand side. This gives
$\sup_{\alpha\in\mathcal{K}}\sup_{P\in\mathcal{P}_{\alpha}}\frac{R_{T}(\hat{\Gamma}_{\bm{u}^{\star}},\hat{\pi}_{\bm{u}^{\star}};P)}{R_{T}^{\star}(\alpha)}\;\lesssim\;M^{5}(\log T)\,T^{\psi_{M}^{\star}}.$
This matches the lower bound of Theorem 1 up to logarithmic factors, completing the proof of Theorem 2.
4.5 Solve the variational problem
The final step in implementing RoBIN is computing the optimal batch grid $\bm{u}^{\star}$, which solves the variational problem (8) underlying our theoretical analysis. Although Proposition 3 establishes that the objective function $\psi(\bm{u})$ is convex in $\bm{u}$, convexity alone does not immediately yield an efficient numerical solver. Instead, we exploit an equivalent characterization of the optimal solution.
Proposition 5.
Define $\phi_{M}(x)=\min_{\alpha\in[0,d/\beta]}\gamma(\alpha)x+h_{M}(\alpha)$ for $x\in(0,1)$. The unique solution $\bm{u}^{\star}$ to the variational problem (8) is also the unique root to the following nonlinear system of equations:
$u_{2}$ $=u_{1}+\phi_{M}(u_{1}),$
$\cdots$
$u_{m}$ $=u_{1}+\phi_{M}(u_{m-1}),$
$\cdots$
$1=u_{M}$ $=u_{1}+\phi_{M}(u_{M-1}).$
See Appendix B for the proof of this proposition.
This equivalence provides a simple and robust computational procedure. Two structural properties are immediate: (1) the final value $u_{M}$ is a strictly increasing function of $u_{1}$, and (2) evaluating the univariate function $\phi_{M}(x)$ amounts to solving the convex problem $\min_{\alpha\in[0,d/\beta]}\gamma(\alpha)x+h_{M}(\alpha)$. Consequently, we can recover $\bm{u}^{\star}$ by a one-dimensional bisection search on $u_{1}$: start with an interval $a,b]\subset(0,1)$ such that $u_{M}(a)<1<u_{M}(b)$; iteratively update $u_{1}$ by halving the interval until $u_{M}$ computed from the recursive relations above equals $1$ within numerical tolerance. This routine yields the optimal grid $\bm{u}^{\star}$ efficiently and stably even for large $M$.
5 Proof of the lower bound
We now establish the lower bound in Theorem 1, proving that every $M$-batch policy must suffer regret inflation of at least order $T^{\psi_{M}^{\star}}$.
5.1 Proof overview
Let $\bm{u}^{\star}$ be a minimizer of the variational problem (8). By Proposition 3(ii), there exist margin parameters $\alpha_{1}=\infty\geq\alpha_{2}\geq\cdots\geq\alpha_{M}$ with $\alpha_{i}\leq d/\beta$ for $i\geq 2$ such that
$\psi_{M}^{\star}\;=\;\eta_{i}(\bm{u}^{\star},\alpha_{i})-h_{M}(\alpha_{i})\quad\text{for each }i\in[M].$ (14)
Define batch cutoffs $T_{0}=0$, $T_{i}=\lceil T^{u_{i}^{\star}}\rceil$ for $1\leq i\leq M-1$, and $T_{M}=T$. We interpret index $i$ as a difficulty level: level 1 has a constant gap (easy), while levels $i\geq 2$ have nonincreasing margin parameters (harder). Crucially, resolving level $i$ requires on the order of $T^{u_{i-1}^{\star}}$ samples, which is exactly the size of the preceding window.
The key insight is that an algorithm that does not know which level it is facing must allocate grids sub-optimally for at least one level. The proof proceeds in three main steps:
  • Hard-instance construction (Section 5.2). We partition the covariate space into $M$ “stripes”, each operating at a different resolution. Within stripe $i$, we randomly place “active cells” with reward gaps aligned with margin $\alpha_{i}$. An algorithm cannot determine which cells are active without sufficient exploration.
  • Regret in the pivotal window (Section 5.3). Let $\Gamma=\{0<t_{1}<t_{2}<\cdots<t_{M-1}<T\}$ be the (possibly adaptively chosen) grid points by the policy. Define the bad events
    $A_{i}=\{t_{i-1}<T_{i-1}<T_{i}\leq t_{i}\},\qquad i=1,\ldots,M,$ (15)
    which form a partition of the sample space. Roughly speaking, $A_{i}$ represents the event that the algorithm’s chosen grid points are suboptimal between the (i1)(i-1)-th and $i$-th batch. We show that if $A_{i}$ happens, the algorithm incurs large regret on level-$i$ cells within the window $T_{i-1},T_{i}]$.
  • Indistinguishability across consecutive levels (Section 5.4). We define mixtures $Q_{i}$ by randomizing the active sets and signs for levels $m\leq i$. The key is that observations up to time $T_{i-1}$ cannot reliably distinguish whether the environment is drawn from $Q_{i}$ or $Q_{i+1}$:
[…]
    This is proved via an equivalent coin model (biased vs. fair coins) and a $\chi^{2}$ bound (Lemma 5), followed by Pinsker. By a telescoping argument over $i$ and the triangle inequality, there exists an index $i^{\star}$ such that
[…]
    Thus, on at least one level, the pivotal window occurs with nontrivial probability under the corresponding mixture.
Figure 3: Visualization of the active cells when $d=2,M=3$. The domain is partitioned into $M=3$ vertical stripes, each subdivided into fine grids of micro-cells. Colored squares indicate active regions, with each color corresponding to a different stripe resolution parameter $z_{m}$.
5.2 The hard instance construction
We now construct the family of hard instances $\{Q_{1},Q_{2},\ldots,Q_{M}\}$.
Step 1: designing the covariate distribution.
Recall that an instance $P$ dictates a law over $X_{t},Y^{(1)}_{t},Y^{(-1)}_{t}$. We begin with describing the covariate distribution $P_{X}$, which is shared among $\{Q_{1},Q_{2},\ldots,Q_{M}\}$.
We start with partition the covariate space $0,1]^{d}$. Define
$z_{1}=1,\quad z_{m}=\lceil 16M^{-1}(M^{5}T^{u^{\star}_{m-1}})^{\frac{1}{2\beta+d}}\rceil,\text{ for }2\leq m\leq M.$
Split coordinate $x_{1}$ into $M$ stripes $\mathcal{S}_{m}=\{x\in[0,1]^{d}:x_{1}\in[(m-1)/M,m/M)\}$, $m=1,\ldots,M$. Fix integers $z_{m}$ and set
$w_{m}=\frac{1}{Mz_{m}},\qquad r_{m}=\frac{1}{4Mz_{m}}.$
Inside stripe $m$, form an axis-aligned grid of micro-cells $\{C_{m,j}\}_{j=1}^{Z_{m}}$ of side-length $w_{m}$ by using $z_{m}$ cuts along $x_{1}$ (within the stripe) and $Mz_{m}$ cuts along each of the remaining d1d-1 coordinates. Thus
$Z_{m}=z_{m}\cdot(Mz_{m})^{d-1}=M^{d-1}z_{m}^{d}.$
Let $q_{m,j}$ be the center of $C_{m,j}$, and define $\ell_{\infty}$ balls
$B_{m,j}\coloneqq B_{\infty}(q_{m,j},r_{m})\subset C_{m,j}.$
With this partition, we define $P_{X}$ to be the uniform distribution on $\bigcup_{m=1}^{M}\bigcup_{j=1}^{Z_{m}}B_{m,j}$. Then
$P_{X}(B_{m,j})=(Mz_{m})^{-d}\qquad\text{for all }m,j.$ (16)
It is straightforward to check that $P_{X}$ obeys Assumption 1.
Step 2: designing the reward family.
Now we are ready to construct the reward functions. Across the families, we will let $f_{(-1)}\equiv\frac{1}{2}$. Fix a bump $\phi:[0,\infty)\to[0,1]$:
$\phi(r)=\begin{cases}1,&0\leq r<\tfrac{1}{4},\\ 2-4r,&\tfrac{1}{4}\leq r<\tfrac{1}{2},\\ 0,&r\geq\tfrac{1}{2}.\end{cases}$
For level $m$, define
$\xi_{m,j}(x)\coloneqq\delta_{m}\,\phi^{\beta}\big(Mz_{m}\|x-q_{m,j}\|_{\infty}\big)\,\mathbf{1}{\{x\in C_{m,j}\}},\qquad\delta_{m}\coloneqq D_{\phi}\,(Mz_{m})^{-\beta},$ (17)
with $D_{\phi}=\min(4^{-\beta}L,1/4)$. Then $\xi_{m,j}$ is supported on $B_{m,j}$, equals $\delta_{m}$ on the inner quarter, and is $(\beta,L)$-Hölder.
Choose a subset $S_{m}\subset[Z_{m}]$ with size
$|S_{m}|=s_{m}\coloneqq\Big\lceil M^{-1}(Mz_{m})^{\,d-\alpha_{m}\beta}\Big\rceil,\quad 2\leq m\leq M,\quad|S_{1}|=s_{1}\coloneqq M^{d-1},$ (18)
and attach i.i.d. Rademacher signs $\{\sigma_{m,i}\}_{i\in S_{m}}$. The subset $S_{m}$ controls which micro-cells are active for the reward function in the $m$-th stripe; see Figure 3 for an illustration.
For $i\in[M]$, define the level-$i$ reward family $\mathcal{F}_{i}$ to be
$\mathcal{F}_{i}=\Big\{\,(f^{(1)}_{S,\sigma,i}(x)=\tfrac{1}{2}+\sum_{m=1}^{i}\ \sum_{j\in S_{m}}\sigma_{m,j}\,\xi_{m,j}(x),\quad f^{(-1)}\equiv\tfrac{1}{2}):\textrm{for all possible configurations of }S,\sigma\,\Big\}.$ (19)
Proposition 6.
For every $i$ and $f^{(1)}_{S,\sigma,i}\in\mathcal{F}_{i}$, the pair $(f^{(1)}_{S,\sigma,i},f^{(-1)})$ is $(\beta,L)$-Hölder and satisfies the margin with parameter $\alpha_{i}$. Hence $\mathcal{F}_{i}\subset\mathcal{P}_{\alpha_{i}}$.
See Section C.1 for the proof.
Step 3: designing the mixture $Q_{i}$.
For $i\in[M]$, define the mixture $Q_{i}$ by drawing $S_{m}$ uniformly among subsets of size $s_{m}$ and i.i.d. signs for $m\leq i$.
5.3 Lower bounding the regret on $Q_{i}$ via indistinguishability
Fix $i\in\{1,\dots,M\}$ and recall $T_{i-1}=\lceil T^{u^{\star}_{i-1}}\rceil$, $T_{i}=\lceil T^{u^{\star}_{i}}\rceil$ with $T_{M}=T$. Let $Q_{i}$ be the level-$i$ mixture from (19) (random $S_{m},\sigma_{m}$ for $m\leq i$) and recall the bad event $A_{i}=\{t_{i-1}<T_{i-1}<T_{i}\leq t_{i}\}$; the $A_{i}$’s partition the sample space. Our goal is to show that, on some $i^{\star}$ for which $Q_{i^{\star}}(A_{i^{\star}})$ is bounded below, the expected regret incurred between rounds $T_{i^{\star}-1}\!+\!1$ and $T_{i^{\star}}$ is large.
Step 1: Restricting to the pivotal window.
Since the maximum is larger than the average and the single-step regret is nonnegative, for any policy $(\Gamma,\pi)$, we have
$\sup_{P\in\mathcal{P}_{\alpha_{i}}}R_{T}(\Gamma,\pi;P)$ $\geq\mathbb{E}_{\mathcal{P}\sim Q_{i}}[R_{T}(\Gamma,\pi;P)]$
$=\mathbb{E}_{\mathcal{P}\sim Q_{i}}\left[\mathbb{E}_{P}\left[\sum_{t=1}^{T}\left(f^{\star}(X_{t})-f^{(\pi_{t}(X_{t}))}(X_{t})\right)\right]\right]$
$\geq\sum_{t=T_{i-1}+1}^{T_{i}}\mathbb{E}_{\mathcal{P}\sim Q_{i}}\left[\mathbb{E}_{P}\left[\left(f^{\star}(X_{t})-f^{(\pi_{t}(X_{t}))}(X_{t})\right)\right]\right].$
Step 2: Localizing to active level-$i$ cells.
For each $1\leq m\leq i$, let $S_{m}$ and $\sigma_{m}$ be randomly and uniformly generated, i.e., $S_{m}$ is a random subset of $Z_{m}]$ with size $s_{m}$, and $\sigma_{m}$ be a random binary vector. Use the definition of $Q_{i}$ to rewrite
$\mathbb{E}_{\mathcal{P}\sim Q_{i}}\left[\mathbb{E}_{P}\left[\left(f^{\star}(X_{t})-f^{(\pi_{t}(X_{t}))}(X_{t})\right)\right]\right]$ $=\mathbb{E}_{\{S_{m}\}_{1\leq m\leq i}}\mathbb{E}_{\{\sigma_{m}\}_{1\leq m\leq i}}\mathbb{E}_{\mathcal{P}_{S,\sigma}}\left[\left(f^{\star}(X_{t})-f^{(\pi_{t}(X_{t}))}(X_{t})\right)\right]$
$=\mathbb{E}_{\{S_{m}\}_{1\leq m\leq i-1}}\mathbb{E}_{\{\sigma_{m}\}_{1\leq m\leq i-1}}\mathbb{E}_{S_{i}}\mathbb{E}_{\sigma_{i}}\mathbb{E}_{\mathcal{P}_{S,\sigma}}\left[\left(f^{\star}(X_{t})-f^{(\pi_{t}(X_{t}))}(X_{t})\right)\right].$
On level $i$ the only locations where the two arms differ are the active micro-cells $\{C_{i,j}\}_{j\in S_{i}}$, and there the gap equals $\delta_{i}$ with sign $\sigma_{i,j}\in\{\pm 1\}$ (see (31) and (32)). Therefore,
$\mathbb{E}_{\mathcal{P}_{S,\sigma}}\left[\left(f^{\star}(X_{t})-f^{(\pi_{t}(X_{t}))}(X_{t})\right)\right]\geq\delta_{i}\cdot\mathbb{E}_{\mathcal{P}_{S,\sigma}}\left[\sum_{j\in S_{i}}\mathbf{1}\{X_{t}\in C_{i,j},\pi_{t}(X_{t})\neq\sigma_{i,j}\}\right]$
Now fix the realization for $\{S_{m}\}_{1\leq m\leq i}$, $\{\sigma_{m}\}_{1\leq m\leq i-1}$, and fix any $j\in S_{i}$. We aim to lower bound $\mathbb{E}_{\sigma_{i}}\mathbb{E}_{\mathcal{P}_{S,\sigma}}\left[\mathbf{1}\{X_{t}\in C_{i,j},\pi_{t}(X_{t})\neq\sigma_{i,j}\}\right]$. Denote by $\sigma_{i,-j}$ the random vector excluding the $j$-th coordinate. We have
$\mathbb{E}_{\sigma_{i}}\mathbb{E}_{\mathcal{P}_{S,\sigma}}\left[\mathbf{1}\{X_{t}\in C_{i,j},\pi_{t}(X_{t})\neq\sigma_{i,j}\}\right]$
$\quad=\frac{1}{2}\mathbb{E}_{\sigma_{i,-j}}\left[\mathbb{P}_{S,\sigma\mid\sigma_{i,j}=1}(X_{t}\in C_{i,j},\pi_{t}(X_{t})\neq 1)+\mathbb{P}_{S,\sigma\mid\sigma_{i,j}=-1}(X_{t}\in C_{i,j},\pi_{t}(X_{t})\neq-1)\right]$
$\quad=\frac{1}{2(Mz_{i})^{d}}\mathbb{E}_{\sigma_{i,-j}}\Big[\underbrace{\mathbb{P}_{S,\sigma\mid\sigma_{i,j}=1}(\pi_{t}(X_{t})\neq 1\mid X_{t}\in C_{i,j})+\mathbb{P}_{S,\sigma\mid\sigma_{i,j}=-1}(\pi_{t}(X_{t})\neq-1\mid X_{t}\in C_{i,j})}_{U^{t}_{i,j}}\Big],$
where we have used the fact that $P_{X}(X_{t}\in C_{i,j})=1/(Mz_{i})^{d}$.
Step 3: Localizing the TV to $A_{i}$ (Le Cam on a subset).
Define $\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}}^{t}$ to be the law of observations up to time $t$ under the environment with $\sigma_{i,j}$ and under the policy $(\Gamma,\pi)$. By Le Cam’s method, one has
$U_{i,j}^{t}$ $\geq 1-\|\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{t}-\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=1}^{t}\|_{\mathrm{TV}}$
$\geq 1-\|\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{T_{i}}-\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=1}^{T_{i}}\|_{\mathrm{TV}}$
$=\int\min\left\{\mathrm{d}\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{T_{i}},\mathrm{d}\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=1}^{T_{i}}\right\}$
$\geq\int_{A_{i}}\min\left\{\mathrm{d}\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{T_{i}},\mathrm{d}\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=1}^{T_{i}}\right\},$
where the second inequality holds since $t\leq T_{i}$. Here we recall that $A_{i}=\{t_{i-1}<T_{i-1}<T_{i}\leq t_{i}\}$. Under $A_{i}$, the available observations at $T_{i}$ are the same as those at $T_{i-1}$ under $A_{i}$, we therefore have
$U_{i,j}^{t}$ $\geq\int_{A_{i}}\min\left\{\mathrm{d}\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{T_{i-1}},\mathrm{d}\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=1}^{T_{i-1}}\right\}$
$=\frac{1}{2}\int_{A_{i}}\left(\mathrm{d}\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{T_{i-1}}+\mathrm{d}\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=1}^{T_{i-1}}-|\mathrm{d}\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{T_{i-1}}-\mathrm{d}\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=1}^{T_{i-1}}|\right)$
$\geq\frac{1}{2}\left(\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{T_{i-1}}(A_{i})+\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=1}^{T_{i-1}}(A_{i})\right)-\|\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{T_{i-1}}-\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=1}^{T_{i-1}}\|_{\mathrm{TV}}.$
For the TV distance, we have the following bound, whose proof is deferred to Section C.2.
Lemma 2.
Fix any $n\in[T]$ and any policy $(\Gamma,\pi)$. For any $i\in[M]$ and $j\in S_{i}$,
$\|\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{n}-\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=1}^{n}\|_{\mathrm{TV}}\leq\sqrt{n(Mz_{i})^{-(2\beta+d)}}.$
Applying Lemma 2 with $n=T_{i-1}$, we obtain
$U_{i,j}^{t}\geq\frac{1}{2}\left(\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{T_{i-1}}(A_{i})+\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=1}^{T_{i-1}}(A_{i})\right)-\frac{1}{4M}.$
Step 4: Averaging over $(S_{m},\sigma_{m})$.
Combining Steps 1-3, we arrive at
$\sup_{P\in\mathcal{P}_{\alpha_{i}}}R_{T}(\Gamma,\pi;P)$
$\geq\frac{\delta_{i}}{2(Mz_{i})^{d}}\sum_{t=T_{i-1}+1}^{T_{i}}\mathbb{E}_{\{S_{m}\}_{1\leq m\leq i-1}}\mathbb{E}_{\{\sigma_{m}\}_{1\leq m\leq i-1}}\mathbb{E}_{S_{i}}\sum_{j\in S_{i}}\mathbb{E}_{\sigma_{i,-j}}\left[\frac{1}{2}\left(\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{T_{i-1}}(A_{i})+\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=1}^{T_{i-1}}(A_{i})\right)-\frac{1}{4M}\right]$
$=\frac{\delta_{i}\cdot s_{i}}{2(Mz_{i})^{d}}\sum_{t=T_{i-1}+1}^{T_{i}}\left(\mathbb{P}_{Q_{i}}^{T_{i-1}}(A_{i})-\frac{1}{4M}\right)$
$=\frac{\delta_{i}\cdot s_{i}}{2(Mz_{i})^{d}}(T_{i}-T_{i-1})\left(\mathbb{P}_{Q_{i}}^{T_{i-1}}(A_{i})-\frac{1}{4M}\right).$
Here the first equality essentially uses the definition of $Q_{i}$.
Since the event $A_{i}$ can be determined by observations up to $T_{i-1}$, we have $\mathbb{P}_{Q_{i}}^{T_{i-1}}(A_{i})=\mathbb{P}_{Q_{i}}(A_{i})$. It boils down to lower bounding $\mathbb{P}_{Q_{i}}(A_{i})$, for which we have the following lemma.
Lemma 3.
There exists some $1\leq i^{\star}\leq M$ such that $\mathbb{P}_{Q_{i^{\star}}}(A_{i^{\star}})\geq 1/(2M)$.
From now on, we identify $i$ with $i^{\star}$. As a result, we have
$\sup_{P\in\mathcal{P}_{\alpha_{i}}}R_{T}(\Gamma,\pi;P)\gtrsim\frac{\delta_{i}\cdot s_{i}}{2(Mz_{i})^{d}}\frac{T_{i}}{M}.$
When $i=1$, $s_{1}=M^{d-1}$, one has
$\sup_{P\in\mathcal{P}_{\alpha_{1}}}R_{T}(\Gamma,\pi;P)\gtrsim\frac{\delta_{1}\cdot s_{1}}{2(Mz_{1})^{d}}\frac{T_{1}}{M}\apprge M^{-3}\cdot T^{u_{1}^{\star}}.$
When $i\geq 2$, recall that $\delta_{i}=D_{\phi}(Mz_{i})^{-\beta}$, $s_{i}=M^{-1}(Mz_{i})^{d-\alpha_{i}\beta}$, $T_{i}\asymp T^{u_{i}^{\star}}$, and $Mz_{i}\asymp(M^{5}T^{u^{\star}_{m-1}})^{\frac{1}{2\beta+d}}$. We therefore obtain
$\sup_{P\in\mathcal{P}_{\alpha_{i}}}R_{T}(\Gamma,\pi;P)\gtrsim\frac{1}{M^{7}}\cdot T^{u_{i}^{\star}-u_{i-1}^{\star}\gamma(\alpha_{i})}.$
Combining the above relations with the inequality
$\mathrm{RI}(\Gamma,\pi)=\sup_{\alpha\in\mathcal{K}}\sup_{P\in\mathcal{P}_{\alpha}}\frac{R_{T}(\Gamma,\pi;P)}{R_{T}^{\star}(\alpha)}\geq\sup_{P\in\mathcal{P}_{\alpha_{i}}}\frac{R_{T}(\Gamma,\pi;P)}{R_{T}^{\star}(\alpha_{i})}$ (20)
yields
$\mathrm{RI}(\Gamma,\pi)\geq\sup_{P\in\mathcal{P}_{\alpha_{i}}}\frac{R_{T}(\Gamma,\pi;P)}{R_{T}^{\star}(\alpha_{i})}\apprge\frac{1}{\log T}\cdot M^{-8}\cdot T^{u_{i}^{\star}-u_{i-1}^{\star}\gamma(\alpha_{i})-h_{M}(\alpha_{i})}\asymp\frac{1}{\log T}\cdot M^{-8}\cdot T^{\psi_{M}^{\star}}.$ (21)
5.4 Proving the indistinguishability
In this section, we aim to demonstrate that the family $\{Q_{1},Q_{2},\ldots,Q_{M}\}$ is indistinguishable from finite samples. As a consequence, we establish Lemma 3.
The following lemma is the key result of this section, which establishes the fact that for any policy $(\Gamma,\pi)$, given observations up to time $T_{i-1}$, it is not possible to distinguish if the bandit instance is from $Q_{i}$ or from $Q_{i+1}$.
Lemma 4.
Fix any policy $(\Gamma,\pi)$. Denote by $Q_{i}^{T_{i-1}}$ the law of observation up to time $T_{i-1}$ under the mixture distribution $Q_{i}$ and under the policy $(\Gamma,\pi)$. Then for any $1\leq i\leq M-1$, one has
$\mathrm{TV}(Q_{i}^{T_{i-1}},Q_{i+1}^{T_{i-1}})\leq\frac{1}{2M^{2}}\cdot T^{u_{i-1}^{\star}-u_{i}^{\star}\cdot(\gamma_{i+1}^{\star}+\frac{1}{2})}.$
Before diving into the proof of this lemma, we prove Lemma 3 based on Lemma 4. By the triangle inequality, we have
$\mathrm{TV}(Q_{i}^{T_{i-1}},Q_{M}^{T_{i-1}})\leq\sum_{m=i}^{M-1}\mathrm{TV}(Q_{m}^{T_{i-1}},Q_{m+1}^{T_{i-1}})$ $\overset{\mathrm{(i)}}{\leq}\sum_{m=i}^{M-1}\mathrm{TV}(Q_{m}^{T_{m-1}},Q_{m+1}^{T_{m-1}})$
$\overset{\mathrm{(ii)}}{\leq}\frac{1}{2M^{2}}\sum_{m=i}^{M-1}T^{u^{\star}_{m-1}-u_{m}^{\star}\cdot(\gamma_{m+1}^{\star}+\frac{1}{2})}$
$\overset{\mathrm{(iii)}}{\leq}\frac{1}{2M}.$ (22)
Here, step (i) uses the fact that $T_{m-1}\geq T_{i-1}$, step (ii) uses Lemma 4, and step (iii) uses the fact that $u^{\star}_{m-1}-u_{m}^{\star}\cdot(\gamma_{m+1}^{\star}+\frac{1}{2})\leq 0$ from Lemma 13.
As a result, we obtain
$|Q_{M}(A_{i})-Q_{i}(A_{i})|\overset{\mathrm{}}{=}|Q_{M}^{T_{i-1}}(A_{i})-Q_{i}^{T_{i-1}}(A_{i})|\overset{\mathrm{}}{\leq}\mathrm{TV}(Q_{M}^{T_{i-1}},Q_{i}^{T_{i-1}})\overset{\mathrm{}}{\leq}\frac{1}{2M},$ (23)
where the first step holds since $A_{i}$ can be determined by observations up to $T_{i-1}$, the second step uses the definition of TV, and the last step is due to relation (5.4).
Consequently,
$\sum_{i=1}^{M}Q_{i}(A_{i})$ $=Q_{M}(A_{M})+\sum_{i=1}^{M-1}Q_{i}(A_{i})$
$=Q_{M}(A_{M})+\sum_{i=1}^{M-1}(Q_{i}(A_{i})-Q_{M}(A_{i})+Q_{M}(A_{i}))$
$\overset{\mathrm{(iv)}}{\geq}Q_{M}(A_{M})+\sum_{i=1}^{M-1}(Q_{M}(A_{i})-\frac{1}{2M})\geq\sum_{i=1}^{M}Q_{M}(A_{i})-\frac{1}{2}\overset{\mathrm{(v)}}{=}\frac{1}{2},$
where step (iv) uses inequality (23), and step (v) uses the fact that $\sum_{i=k}^{M}Q_{M}(A_{i})=1$. Lemma 3 follows from the pigeonhole principle.
Now we return to the proof of Lemma 4. When $i=1$, one has $T_{0}=0$ and the statement trivially holds. Hence in the remaining proof, we consider $i\geq 2$.
5.4.1 An equivalent coin model
In this section, we introduce an equivalent coin model to help us control $\mathrm{TV}(Q_{i}^{T_{i-1}},Q_{i+1}^{T_{i-1}})$. The coin model is indexed by four parameters: $z$, the number of coins, $s$, the number of possibly biased coins, $\delta$, the effective bias of the coin, and $n$, the total number of tosses.
Suppose there are $z$ coins labeled by $1,2,\ldots,z$. We perform $n$ rounds of experiments. In each round $t$: we pick a random coin $I_{t}\sim\text{Unif}\{1,\ldots,z\}$ independently, flip that coin, and observe the outcome $Y_{t}\in\{0,1\}$. We define for each coin $i$:
$N_{i}\coloneqq\sum_{t=1}^{N}\mathbf{1}\{I_{t}=i\},\quad R_{i}\coloneqq\sum_{t=1}^{N}\mathbf{1}\{I_{t}=i,Y_{t}=1\}.$
In words, $N_{i}$ is the number of tosses for coin $i$, and $R_{i}$ is the number of heads for coin $i$.
We consider two possible hypothesis for the bias of the coins.
Null model $H_{0}$.
Under $H_{0}$, every coin is fair: the probability of heads $p_{i}=1/2$ for all $i$. Conditional on the number of times a coin was used:
$R_{i}\mid N_{i}\sim\text{Bin}(N_{i},1/2),$
and the different coins’ results $(R_{i})_{i=1}^{z}$ are independent given the counts $(N_{i})_{i=1}^{z}$. Denote by $P_{0}$ the joint law of the observed data under $H_{0}$.
Alternative model $H_{1}$.
Now, under $H_{1}$, we introduce a small number of biased coins by randomly choosing a subset of coins
$S\subset[z],\quad|S|=s,$
uniformly among all $s$-element subsets. Also, Let $\sigma_{i}\in\{\pm 1\}$ be i.i.d. Rademacher random variables for $i\in[z]$. For coins $i\in S$, they are biased either upwards or downwards,
$R_{i}\mid(S,\sigma,N_{i})\sim\text{Bin}\left(N_{i},\frac{1}{2}+\sigma_{i}\delta\right),$
where $\delta\in(0,1/2)$ is the bias magnitude. Coins $i\notin S$ are still fair,
$R_{i}\mid(S,\sigma,N_{i})\sim\text{Bin}(N_{i},1/2).$
Denote by $P_{S,\sigma}$ the joint law of the observed data given $S,\sigma$. Define $Q=\mathbb{E}_{S,\sigma}[P_{S,\sigma}]$ to be the mixture distribution under $H_{1}$.
We have the following control on the chi-squared divergence between the null model and the alternative model.
Lemma 5.
Assume that $n\delta^{2}z^{-1}<o(1)$ and $n^{2}s^{2}\delta^{4}/z^{3}\leq 1/32$, then
$\chi^{2}(Q,P_{0})\leq\frac{32n^{2}s^{2}\delta^{4}}{z^{3}}.$
5.4.2 Connect $\mathrm{TV}(Q_{i}^{T_{i-1}},Q_{i+1}^{T_{i-1}})$ to the coin model
Define $Q_{i}^{\otimes T_{i-1}}$ to be the joint law of the full observations
$(X_{t},Y_{t}^{(1)},Y_{t}^{(-1)}),\quad 1\leq t\leq T_{i-1},$
under the mixture distribution $Q_{i}$. It is worth noting that $Q_{i}^{\otimes T_{i-1}}$ is independent from any policy $(\Gamma,\pi)$, as opposed to $Q_{i}^{T_{i-1}}$.
By the data processing inequality, we know that
$\mathrm{TV}(Q_{i}^{T_{i-1}},Q_{i+1}^{T_{i-1}})\leq\mathrm{TV}(Q_{i}^{\otimes T_{i-1}},Q_{i+1}^{\otimes T_{i-1}}).$ (24)
Recall the definitions of $Q_{i}$ and $Q_{i+1}$. We note that they only differ when $X_{t}\in\mathcal{S}_{i+1}$. Due to the Bernoulli reward structure, in this region, $Q_{i+1}$ now corresponds to the alternative model $H_{1}$ where a subset of coins are biased and $Q_{i}$ corresponds to the null model $H_{0}$. Since only samples landing into $\mathcal{S}_{i+1}$ can help distinguish $Q_{i}^{\otimes T_{i-1}}$ and $Q_{i+1}^{\otimes T_{i-1}}$, one has
$\mathrm{TV}(Q_{i}^{\otimes T_{i-1}},Q_{i+1}^{\otimes T_{i-1}})\leq\mathrm{TV}(P_{0},Q),$ (25)
where $P_{0},Q$ denote the coin model with parameters $z=Z_{i+1}$, $s=s_{i+1}^{\mathrm{tot}}$, $\delta=\delta_{i+1}$, and $n=T_{i-1}$.
Under the choice of $n,\delta,z$ and $s$, it can be verified that $n\delta^{2}z^{-1}<o(1)$ and $n^{2}s^{2}\delta^{4}/z^{3}\leq 1/32$. Hence, by Pinsker’s inequality and Lemma 5,
$\mathrm{TV}(P_{0},Q)\leq\sqrt{\frac{1}{2}\chi^{2}(Q,P_{0})}\leq\sqrt{\frac{16n^{2}s^{2}\delta^{4}}{z^{3}}}=\frac{4ns\delta^{2}}{z^{1.5}}.$ (26)
Combining relations (24), (25) and (26),
$\mathrm{TV}(Q_{i}^{T_{i-1}},Q_{i+1}^{T_{i-1}})\leq\frac{4ns\delta^{2}}{z^{1.5}}=\frac{4T_{i-1}s_{i+1}\delta_{i+1}^{2}}{Z_{i+1}^{1.5}}\leq\frac{1}{2M^{2}}\cdot T^{u_{i-1}^{\star}-u_{i}^{\star}\cdot(\gamma_{i+1}^{\star}+\frac{1}{2})}.$
This completes the proof of Lemma 4. The remaining of this section is devoted to proving Lemma 5.
5.4.3 Proof of Lemma 5
We bound the chi-squared divergence using the second-moment representation, i.e.,
$\chi^{2}(Q,P_{0})=\mathbb{E}_{P_{0}}[\mathbb{E}[\Lambda^{2}\mid\bm{N}]]-1,$
where we further condition on $\bm{N}=(N_{1},\dots,N_{z})$, the multinominal vector which counts the number of times each coin is flipped.
The conditional likelihood ratio.
Conditioned on $\bm{N}$, coins are independent under both $H_{0}$ and $H_{1}$. Denote by $\bm{R}=(R_{1},\dots,R_{z})$. For the null model,
$P_{0}(R\mid\bm{N})=\prod_{i=1}^{z}\binom{N_{i}}{R_{i}}2^{-N_{i}}.$
Under a fixed $(S,\sigma)$,
$P_{S,\sigma}(\bm{R}\mid\bm{N})=\prod_{i\notin S}\binom{N_{i}}{R_{i}}2^{-N_{i}}\prod_{i\in S}\binom{N_{i}}{R_{i}}\left(\frac{1}{2}+\sigma_{i}\delta\right)^{R_{i}}\left(\frac{1}{2}-\sigma_{i}\delta\right)^{N_{i}-R_{i}}.$
Hence the per-coin likelihood ratio factor for $i\in S$ is
$r_{i}(\sigma_{i})=\frac{\left(\frac{1}{2}+\sigma_{i}\delta\right)^{R_{i}}\left(\frac{1}{2}-\sigma_{i}\delta\right)^{N_{i}-R_{i}}}{(1/2)^{N_{i}}}=(1+2\sigma_{i}\delta)^{R_{i}}(1-2\sigma_{i}\delta)^{N_{i}-R_{i}}.$
Consequently,
$Q(\bm{R}\mid\bm{N})=\frac{1}{\binom{z}{s}}\sum_{S:\,|S|=s}\mathbb{E}_{\sigma}\!\left[\,P_{S,\sigma}(\bm{R}\mid\bm{N})\,\right].$
Dividing by $P_{0}(\bm{R}\mid\bm{N})$ gives
$\Lambda(\bm{R}\mid\bm{N})=\frac{1}{\binom{z}{s}}\sum_{S:\,|S|=s}\mathbb{E}_{\sigma}\!\left[\prod_{i\in S}r_{i}(\sigma_{i})\right],$
where we use $\Lambda$ to denote the likelihood ratio between $Q$ and $P_{0}$. Because the $\sigma_{i}$’s are i.i.d., the expectation factorizes:
$\mathbb{E}_{\sigma}\!\left[\prod_{i\in S}r_{i}(\sigma_{i})\right]=\prod_{i\in S}m_{i},\quad\text{where}\quad m_{i}\coloneqq\frac{1}{2}\big(r_{i}(+1)+r_{i}(-1)\big).$
Putting it together,
$\Lambda(\bm{R}\mid\bm{N})=\frac{1}{\binom{z}{s}}\sum_{S:\,|S|=s}\prod_{i\in S}m_{i}.$
Representation of $\Lambda^{2}$.
To control the chi-square divergence, it suffices to bound the second moment of the likelihood ratio. We compute
$\mathbb{E}_{P_{0}}[\Lambda^{2}\mid\bm{N}]$ $=\mathbb{E}_{P_{0}}\left[\left(\frac{1}{\binom{z}{s}}\sum_{S\subset[z],\,|S|=s}\prod_{i\in S}m_{i}\right)\left(\frac{1}{\binom{z}{s}}\sum_{S^{\prime}\subset[z],\,|S^{\prime}|=s}\prod_{j\in S^{\prime}}m_{j}\right)\,\Big|\,\bm{N}\right]$
$=\frac{1}{\binom{z}{s}^{2}}\sum_{\begin{subarray}{c}|S|=s\\ |S^{\prime}|=s\end{subarray}}\mathbb{E}_{P_{0}}\!\left[\prod_{i\in S}m_{i}\prod_{j\in S^{\prime}}m_{j}\,\Big|\,\bm{N}\right].$
Fix any ordered pair $(S,S^{\prime})$. For each index $k\in[z]$, its contribution to the product $\left(\prod_{i\in S}m_{i}\right)\left(\prod_{j\in S^{\prime}}m_{j}\right)$ depends on which of the sets $S,S^{\prime}$ it belongs to:
  • If $k\in S\cap S^{\prime}$: the factor contributes $m_{k}\cdot m_{k}=m_{k}^{2}$.
  • If $k\in S\setminus S^{\prime}$ or $k\in S^{\prime}\setminus S$: the factor contributes a single $m_{k}$.
  • If $k\notin S\cup S^{\prime}$: the factor contributes $1$.
Define the function $g_{b}(a)=((1+4b^{2})^{a}+(1-4b^{2})^{a})/2$ for some $0<b<1/2$. The following lemma helps control the moments of $m$, whose proof is deferred to Section C.3.
Lemma 6.
Under $P_{0}$ and conditional on $\bm{N}$,
$\mathbb{E}_{P_{0}}[m_{i}\mid N_{i}]=1,\quad\mathbb{E}_{P_{0}}[m_{i}^{2}\mid N_{i}]=\tfrac{1}{2}\Big((1+4\delta^{2})^{N_{i}}+(1-4\delta^{2})^{N_{i}}\Big)=g_{\delta}(N_{i}).$
Hence,
$\mathbb{E}_{P_{0}}\!\left[\prod_{i\in S}m_{i}\prod_{j\in S^{\prime}}m_{j}\,\Big|\,\bm{N}\right]=\left(\prod_{i\in S\cap S^{\prime}}\mathbb{E}[m_{i}^{2}]\right)\left(\prod_{i\in S\triangle S^{\prime}}\mathbb{E}[m_{i}]\right)=\prod_{i\in S\cap S^{\prime}}g_{\delta}(N_{i}),$
where $S\triangle S^{\prime}=(S\setminus S^{\prime})\cup(S^{\prime}\setminus S)$ is the symmetric difference.
Averaging over the randomness of $\bm{N}$.
By the law of total expectation, we reach
$\mathbb{E}_{P_{0}}[\mathbb{E}[\Lambda^{2}\mid\bm{N}]]$ $=\frac{1}{\binom{z}{s}^{2}}\sum_{\begin{subarray}{c}|S|=s\\ |S^{\prime}|=s\end{subarray}}\mathbb{E}_{P_{0}}\left[\prod_{i\in S\cap S^{\prime}}g_{\delta}(N_{i})\right]$
$\leq\frac{1}{\binom{z}{s}^{2}}\sum_{\begin{subarray}{c}|S|=s\\ |S^{\prime}|=s\end{subarray}}\prod_{i\in S\cap S^{\prime}}\mathbb{E}_{P_{0}}[g_{\delta}(N_{i})]$
$=\frac{1}{\binom{z}{s}^{2}}\sum_{\begin{subarray}{c}|S|=s\\ |S^{\prime}|=s\end{subarray}}\left[\frac{1}{2}\left((1+4\frac{\delta^{2}}{z})^{n}+(1-4\frac{\delta^{2}}{z})^{n}\right)\right]^{|S\cap S^{\prime}|},$
where the second step is due to the negative association of multinomial random variables (Joag-Dev and Proschan, 1983), and the last step applies the probability generating function of the binomial distribution. We reach
$\mathbb{E}_{P_{0}}[\mathbb{E}[\Lambda^{2}\mid\bm{N}]]$ $=\mathbb{E}_{S,S^{\prime}}\left[(g_{\delta/\sqrt{z}}(n))^{|S\cap S^{\prime}|}.\right]$
We record a useful lemma for controlling the generating function of the average intersection size.
Lemma 7.
Let $S,S^{\prime}$ be indepndent $s$-subsets of $[z]$ and let $L=|S\cap S^{\prime}|$. For any $t\geq 1$,
$\mathbb{E}[t^{L}]\leq\exp(\frac{s^{2}}{z}(t-1)).$
See Section C.3 for the proof.
Applying Lemma 7,
$\mathbb{E}_{P_{0}}[\mathbb{E}[\Lambda^{2}\mid\bm{N}]]$ $=\mathbb{E}_{S,S^{\prime}}\left[(g_{\delta/\sqrt{z}}(n))^{|S\cap S^{\prime}|}\right]\leq\exp\left(\frac{s^{2}}{z}(g_{\delta/\sqrt{z}}(n)-1)\right).$ (27)
Denote by $\epsilon=4\delta^{2}/z$. By definition,
$g_{\delta/\sqrt{z}}(n)$ $=\frac{1}{2}\left((1+\epsilon)^{n}+(1-\epsilon)^{n}\right)$
$\leq\frac{1}{2}\left(\exp(n\epsilon)+\exp(-n\epsilon)\right)$
$=\cosh(n\epsilon)=1+\frac{(n\epsilon)^{2}}{2}+O((n\epsilon)^{4}),$
where the second step is due to the elementary inequality $1+x\leq e^{x}$, and the last step is by the assumption $n\epsilon<o(1)$. Plugging the above back to (27),
$\mathbb{E}_{P_{0}}[\mathbb{E}[\Lambda^{2}\mid\bm{N}]]$ $\leq\exp\left(\frac{s^{2}}{z}(g_{\delta/\sqrt{z}}(n)-1)\right)$
$\leq\exp\left(\frac{s^{2}}{z}\cdot(n\epsilon)^{2}\right)=\exp\left(\frac{16n^{2}s^{2}\delta^{4}}{z^{3}}\right),$
where the second inequality holds for $n\epsilon$ sufficiently small.
Putting things together.
By the definition of chi-squared divergence,
$\chi^{2}(Q,P_{0})$ $=\mathbb{E}_{P_{0}}[\mathbb{E}[\Lambda^{2}\mid\bm{N}]]-1\leq\exp\left(\frac{16n^{2}s^{2}\delta^{4}}{z^{3}}\right)-1\leq\frac{32n^{2}s^{2}\delta^{4}}{z^{3}},$
where the last step is due to the assumption $n^{2}s^{2}\delta^{4}/z^{3}\leq 1/32$ and the elementary inequality $e^{x}\leq 2x+1$ when $0\leq x\leq 1$.
6 Discussion
This work provides a complete characterization of the cost of adaptivity in batched nonparametric contextual bandits. By introducing the notion of regret inflation, we quantify how much additional regret is unavoidable when the margin parameter $\alpha$ is unknown. Our main finding is that this adaptivity cost scales as $T^{\psi_{M}^{\star}}$, where $\psi_{M}^{\star}$ is the value of a convex variational problem depending on the number of batches $M$, the smoothness $\beta$, and the dimension $d$. The matching upper and lower bounds show that this exponent captures the exact statistical price of limited adaptivity, up to logarithmic factors.
6.1 Phase transition in batched contextual bandits
Our results, together with those in Jiang and Ma (2025) (see also Proposition 2 in this paper), reveal a sharp phase transition in the interplay between batching, adaptivity, and oracle knowledge.
Figure 4 provides a clear visual summary of this landscape. The green curve represents the normalized regret exponent of an oracle algorithm (e.g., BaSEDB) that knows $\alpha$, while the red curve shows the regret exponent of the optimal adaptive algorithm (e.g., RoBIN). The gap between them, which we characterize by $\psi_{M}^{\star}$, is precisely the “price of adaptation.” Our analysis explains the behavior of this gap, revealing three qualitatively distinct regimes determined by the number of batches $M$ relative to the horizon $T$:
  1. 1.
    When $M\leq c\log\log T$ for a sufficiently small constant $c$, the batch constraint itself dominates. In this under-batched regime, even an oracle algorithm that knows the true margin parameter $\alpha$ cannot attain the fully online regret rate.
  2. 2.
    In the intermediate range $c\log\log T<M<C\log\log T$, knowing $\alpha$ becomes beneficial. An oracle algorithm with knowledge of the true margin parameter can achieve the online rate up to logarithmic factors, but ignorance of $\alpha$ induces regret inflation quantified by $\psi_{M}^{\star}$. This transition regime therefore marks the true adaptivity barrier, where batching and parameter uncertainty interact nontrivially.
  3. 3.
    Finally, when $M\geq C\log\log T$ for a sufficiently large constant $C$, the regret inflation becomes constant. In this fully adaptive regime, the proposed algorithm RoBIN matches the performance of the online oracle up to polylogarithmic factors, and adaptation to the unknown $\alpha$ becomes effectively free.
Figure 4: Phase diagram illustrating the three regimes of adaptivity under the batch constraint. The $x$-axis represents the number of batches $M$, and the $y$-axis shows normalized regret. The green curve corresponds to the performance of the optimal algorithm that knows the margin parameter $\alpha$, while the red curve corresponds to the optimal adaptive algorithm without knowledge of $\alpha$. The dotted line shows the exponent for an optimal online algorithm.
These three regimes collectively characterize the adaptivity landscape of batched contextual bandits: batching imposes a fundamental statistical cost only below the doubly logarithmic threshold, beyond which adaptation incurs no additional penalty. (More precisely, the lower and upper thresholds satisfy $c\asymp 1/\log(1/\gamma(\alpha))$ and $C\asymp 1/\log(1/\gamma_{\mathrm{max}})$, where $\gamma_{\mathrm{max}}=\gamma(d/\beta)$ and $\gamma(\alpha)\leq\gamma_{\mathrm{max}}$, giving the ordering $c<C$.)
6.2 Extension to general knowledge sets
In practice, one may have partial prior information about the possible range of the margin parameter. This can be represented by restricting the knowledge set $\mathcal{K}$ to a subset of $0,d/\beta]\cup\{\infty\}$, as in Definition 1. Our main results, in particular Theorem 1, continue to hold whenever $\mathcal{K}\subseteq[d/(2\beta),d/\beta]\cup\{\infty\}$.
Theorem 3 (General knowledge set).
Fix $\mathcal{K}$ to be a compact subset of $d/(2\beta),d/\beta]\cup\{\infty\}$. Let $\psi_{M,\mathcal{K}}^{\star}$ denote the optimal value of the variational problem defined in (8) with knowledge set $\mathcal{K}$. Then there exist constants $c_{1},c_{2}>0$, independent of $T$ and $M$, such that
$c_{1}M^{-8}(\log T)^{-1}T^{\psi_{M,\mathcal{K}}^{\star}}\;\leq\;\inf_{\Gamma,\pi}\mathrm{RI}(\Gamma,\pi)\;\leq\;c_{2}\,M^{5}(\log T)\,T^{\psi_{M,\mathcal{K}}^{\star}}.$
The proof follows closely that of Theorem 1, except that Lemma 13 now holds automatically. Indeed, for all $\alpha\in[d/(2\beta),d/\beta]\cup\{\infty\}$, we have $\gamma(\alpha)=\frac{(\alpha+1)\beta}{d+2\beta}\geq 1/2$, which guarantees that the successive hard-instance families $\{Q_{1},\ldots,Q_{M}\}$ cannot be reliably distinguished within a finite number of batches. This property—anchored at the boundary point $d/(2\beta)$—preserves the indistinguishability argument used in the lower bound proof. By contrast, when the knowledge set $\mathcal{K}$ extends below this threshold, $\gamma(\alpha)$ may fall below $1/2$, breaking the reduction and invalidating the argument.
Nevertheless, if one restricts the $M$-batch policy $(\Gamma,\pi)$ to use a fixed grid, that is, grid points $(t_{1},\ldots,t_{M-1})$ are chosen in advance, then $\psi_{M,\mathcal{K}}^{\star}$ still characterizes the optimal regret inflation over this restricted policy class.
Theorem 4 (General knowledge set, fixed grid).
Fix $\mathcal{K}$ to be a compact subset of $0,d/\beta]\cup\{\infty\}$. Then there exist constants $c_{1},c_{2}>0$, independent of $T$ and $M$, such that
$c_{1}M^{-8}(\log T)^{-1}T^{\psi_{M,\mathcal{K}}^{\star}}\;\leq\;\inf_{\Gamma,\pi}\mathrm{RI}(\Gamma,\pi)\;\leq\;c_{2}\,M^{5}(\log T)\,T^{\psi_{M,\mathcal{K}}^{\star}},$
where the infimum is taken over all $M$-batch policies with a fixed grid.
The upper bound follows directly from the performance guarantee of RoBIN. For the lower bound, since the policy class is restricted to fixed grids, the indistinguishability argument is unnecessary—one can directly choose the worst-case family $Q_{i}$ for the bad event $A_{i}$ (see Equation (15)). We omit the detailed proof for brevity.
6.3 Monotonicity of $\psi_{M}^{\star}$ in $M$
Empirically, $\psi_{M}^{\star}$ decreases monotonically with the number of batches $M$ (see Section 3.5). This behavior is not immediate from the definition of the variational objective in (8): while adding batches allows a finer schedule and hence lower regret, the benchmark $h_{M}(\alpha)$ also shrinks, making the improvement subtle to analyze. For small $M$, we can establish monotonicity formally:
Proposition 7.
For $M\leq 4$, we have $\psi_{M+1}^{\star}<\psi_{M}^{\star}$.
The proof appears in Section B.2.3. Whether this monotonicity extends to all $M$ remains an interesting open question.
6.4 Future directions
Several natural extensions remain open. First, our analysis focuses on the two-armed setting; extending these results to multiple arms is a natural next step. Second, when both the margin and smoothness parameters are unknown, one may ask whether a unified adaptive strategy is possible. Smoothness adaptation is known to require additional structure or to incur unavoidable penalties even in the fully online case (Qian and Yang, 2016b; Gur et al., 2022; Cai et al., 2024). Under batch constraints, the challenge intensifies because the size of the initial batch must be determined before smoothness can be estimated. A joint variational formulation over $(\alpha,\beta)$ may shed light on this problem.
Finally, the regret inflation framework may extend to other batched bandit models. For example, in high-dimensional sparse linear contextual bandits (Ren and Zhou, 2023), existing algorithms rely on knowing an upper bound on the sparsity level. Quantifying the price of unknown sparsity or other structural complexity parameters would further broaden the reach of our theory.
Acknowledgements
C.M. was partially supported by the National Science Foundation via grant DMS-2311127 and the CAREER Award DMS-2443867.
Appendix A Proof of oracle regret bounds (Proposition 2)
We split the proof into two cases: $\alpha=\infty$ and $\alpha\leq d/\beta$.
A.1 The case of $\alpha=\infty$
Without loss of generality, let $\mu_{1}=\mu^{*}$ and $\mu_{2}=\mu^{*}-\Delta$ be the mean rewards of two arms, where $\Delta>0$ is a fixed constant. Clearly, one has $R_{T}^{\star}(\alpha)\geq c_{1}$ for some constant $c_{1}>0$.
A simple policy.
To achieve the upper bound, we consider the following procedure. Given any batch budget $M\geq 2$, we choose to use two batches by setting $t_{1}\asymp\log(T)$. Define the confidence radius
$r(t_{1})=\sqrt{\frac{\log(4T/\delta)}{2t_{1}}},$
where $\delta=1/T^{2}$. During the first batch, we pull each arm in a round-robin fashion. At the end of batch 1, we eliminate any arm $i\in\{1,2\}$ such that
$\widehat{\mu}_{i}(t_{1})+r(t_{1})<\max_{j\in\{1,2\}}\big\{\widehat{\mu}_{j}(t_{1})-r(t_{1})\big\}.$
During the second batch, we just pull any active arm.
Regret analysis.
Now we establish the regret guarantee of the above policy. The following lemma ensures that with high probability, the suboptimal arm is eliminated.
Lemma 8.
With probability at least $1-\delta$, the suboptimal arm is eliminated by phase
$t_{1}=\left\lceil\frac{8}{\Delta^{2}}\log\!\left(\frac{4T}{\delta}\right)\right\rceil.$
Proof.
By Hoeffding’s inequality for bounded rewards,
$\Pr\!\big(|\widehat{\mu}_{i}(t_{1})-\mu_{i}|>r(t_{1})\big)\leq\frac{\delta}{2T}.$
Taking a union bound over both arms,
$\mathcal{E}=\Big\{i\in\{1,2\}:\ |\widehat{\mu}_{i}(t_{1})-\mu_{i}|\leq r(t_{1})\Big\}$
holds with probability at least $1-\delta$.
On the event $\mathcal{E}$,
$\widehat{\mu}_{1}(t_{1})-\widehat{\mu}_{2}(t_{1})\geq(\mu_{1}-\mu_{2})-|\widehat{\mu}_{1}(t_{1})-\mu_{1}|-|\widehat{\mu}_{2}(t_{1})-\mu_{2}|\geq\Delta-2r(t_{1}).$
If $r(t_{1})\leq\Delta/4$, then
$\widehat{\mu}_{1}(t_{1})-\widehat{\mu}_{2}(t_{1})\geq\Delta/2>2r(t_{1}),$
which implies
$\widehat{\mu}_{2}(t_{1})+r(t_{1})<\widehat{\mu}_{1}(t_{1})-r(t_{1}),$
so the suboptimal arm (arm 2) is eliminated at phase $t_{1}$.
The condition $r(t_{1})\leq\Delta/4$ means
$\sqrt{\frac{\log(4T/\delta)}{2t_{1}}}\leq\frac{\Delta}{4}\quad\Longleftrightarrow\quad t_{1}\geq\frac{8}{\Delta^{2}}\log\!\left(\frac{4T}{\delta}\right).$
Thus, on event $\mathcal{E}$ (which holds with probability at least $1-\delta$), arm 2 is eliminated by $t_{1}$. ∎
Denote by $\mathcal{G}$ the event that the suboptimal arm is eliminated by $t_{1}$. By Lemma 8, we have
$R_{T}(\bar{\pi},\mathcal{P})\leq t_{1}\cdot\Delta+(T-t_{1})\cdot\Delta\cdot\delta\leq c_{2}t_{1},$
where we have used the fact that during the second batch regret is only incurred when $\mathcal{G}^{c}$ occurs and $\mathbb{P}(\mathcal{G}^{c})\leq\delta=1/T^{2}$.
A.2 The case of $\alpha\leq d/\beta$
For the remaining of the proof, we establish the result for $\alpha\leq d/\beta$, which is stated in the following proposition.
Proposition 8.
Suppose that $\alpha\leq d/\beta$. Under Assumptions 1-3. For any $M$-batch policy $(\Gamma,\pi)$, one has
$M^{-4}\cdot T^{h_{M}(\alpha)}\apprle\sup_{\mathcal{P}\in\mathcal{P}_{\alpha}}R_{T}(\pi,\mathcal{P})\apprle M(\log T)\cdot T^{\,h_{M}(\alpha)}.$
We use the remaining of the section to prove the above proposition.
A.3 Proof of the upper bound
Define
$g_{0}=\lfloor b^{\frac{1}{2\beta+d}}\rfloor,\qquad\text{and}\qquad g_{i}=\lfloor g_{i-1}^{\gamma}\rfloor,i=1,...,M-2.$ (28)
Denote by $w_{i}=(\prod_{l=0}^{i-1}g_{l})^{-1}$. In addition, define
$t_{i}-t_{i-1}$ $=\lfloor l_{i}w_{i}^{-(2\beta+d)}\log(Tw_{i}^{d})\rfloor,1\leq i\leq M-1,$ (29)
[…]
$R_{T}(\hat{\Gamma},\hat{\pi};P)$ $\apprle t_{1}+\sum_{i=2}^{M-1}(t_{i}-t_{i-1})\cdot w_{i-1}^{\beta+\alpha\beta}+(T-t_{M-1})w_{M-1}^{\beta+\alpha\beta}.$
Under the choices for the batch size and the split factors in (29)-(28),
$t_{1}$ $\lesssim T^{\frac{1-\gamma}{1-\gamma^{M}}}\log T,$
$(t_{i}-t_{i-1})\cdot w_{i-1}^{\beta+\alpha\beta}$ $\lesssim T^{\frac{1-\gamma}{1-\gamma^{M}}}\log T,\qquad\text{for }2\leq i\leq M-1,$
$(T-t_{M-1})w_{M-1}^{\beta+\alpha\beta}$ $\leq Tw_{M-1}^{\beta+\alpha\beta}\lesssim T^{\frac{1-\gamma}{1-\gamma^{M}}}\log T.$
Combining the above three bounds completes the proof.
A.4 Proof of the lower bound
The proof mainly follows the strategy outlined in Jiang and Ma (2025), but with a slightly different reward function construction to handle the wider range of $\alpha$.
A.4.1 Construction of the hard instances
Define $b\asymp T^{(1-\gamma)/(1-\gamma^{M})}$. For each $1\leq m\leq M$, we set $T_{m}=\lfloor b^{(1-\gamma^{m})/(1-\gamma)}\rfloor$. Besides, define
$z_{1}=1,\quad z_{m}=\lceil M^{-1}(36T_{m-1}M^{2})^{1/(2\beta+d)}\rceil,\quad\textrm{for }2\leq m\leq M.$
Constructing the covariate distribution.
Split coordinate $x_{1}$ into $M$ stripes $\mathcal{S}_{m}=\{x\in[0,1]^{d}:x_{1}\in[(m-1)/M,m/M)\}$, $m=1,\ldots,M$. Fix integers $z_{m}$ and set
$w_{m}=\frac{1}{Mz_{m}},\qquad r_{m}=\frac{1}{4Mz_{m}}.$
Inside stripe $m$, form an axis-aligned grid of micro-cells $\{C_{m,j}\}_{j=1}^{Z_{m}}$ of side-length $w_{m}$ by using $z_{m}$ cuts along $x_{1}$ (within the stripe) and $Mz_{m}$ cuts along each of the remaining d1d-1 coordinates. Thus
$Z_{m}=z_{m}\cdot(Mz_{m})^{d-1}=M^{d-1}z_{m}^{d}.$
Let $q_{m,j}$ be the center of $C_{m,j}$, and define $\ell_{\infty}$ balls
$B_{m,j}\coloneqq B_{\infty}(q_{m,j},r_{m})\subset C_{m,j}.$
With this partition, we define $P_{X}$ to be the uniform distribution on $\bigcup_{m=1}^{M}\bigcup_{j=1}^{Z_{m}}B_{m,j}$. Then
$P_{X}(B_{m,j})=(Mz_{m})^{-d}\qquad\text{for all }m,j.$ (30)
It is straightforward to check that $P_{X}$ obeys Assumption 1.
Designing the reward family.
Now we are ready to construct the reward functions. Across the families, we will let $f_{(-1)}\equiv\frac{1}{2}$. Fix a bump $\phi:[0,\infty)\to[0,1]$:
$\phi(r)=\begin{cases}1,&0\leq r<\tfrac{1}{4},\\ 2-4r,&\tfrac{1}{4}\leq r<\tfrac{1}{2},\\ 0,&r\geq\tfrac{1}{2}.\end{cases}$
For level $m$, define
$\xi_{m,j}(x)\coloneqq\delta_{m}\,\phi^{\beta}\big(Mz_{m}\|x-q_{m,j}\|_{\infty}\big)\,\mathbf{1}{\{x\in C_{m,j}\}},\qquad\delta_{m}\coloneqq D_{\phi}\,(Mz_{m})^{-\beta},$ (31)
with $D_{\phi}=\min(4^{-\beta}L,1/4)$. Then $\xi_{m,j}$ is supported on $B_{m,j}$, equals $\delta_{m}$ on the inner quarter, and is $(\beta,L)$-Hölder.
Choose a subset $S_{m}\subset[Z_{m}]$ with size
$|S_{m}|=s_{m}\coloneqq\Big\lceil M^{-1}(Mz_{m})^{\,d-\alpha\beta}\Big\rceil,\quad 2\leq m\leq M,\quad|S_{1}|=s_{1}\coloneqq M^{d-1}.$ (32)
Let $Z=\sum_{m=1}^{M}Z_{m}$. Denote by $\Omega=\{\pm 1\}^{Z}$. We define the reward family $\mathcal{F}$ to be
$\mathcal{F}=\Big\{\,(f^{(1)}_{\omega}(x)=\tfrac{1}{2}+\sum_{m=1}^{M}\ \sum_{j\in S_{m}}\omega_{m,j}\,\xi_{m,j}(x),\quad f^{(-1)}\equiv\tfrac{1}{2}):\omega\in\Omega\,\Big\}.$ (33)
By Proposition 6, we have $\mathcal{F}\subset\mathcal{P}_{\alpha}$.
A.4.2 Lower bounding the regret during the $m$-th batch
Since the worst-case regret is lower bounded by the average regret over the family $\Omega$,
$\sup_{(f,\frac{1}{2})\in\mathcal{F}}$ $R_{T}(\pi,f)$
$\geq\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{E}_{\pi,\omega}\left[\sum_{t=1}^{T}\left(f^{\star}(X_{t})-f^{(\pi_{t}(X_{t}))}(X_{t})\right)\right]$
$\overset{\mathrm{(i)}}{\geq}\sum_{t=T_{m-1}+1}^{T_{m}}\sum_{j\in S_{m}}\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{E}_{\pi,\omega}^{t}\left[D_{\phi}(Mz_{m})^{-\beta}\mathbf{1}\{X_{t}\in B_{m,j},\pi_{t}(X_{t})\neq\omega_{m,j}\}\right]$
$=D_{\phi}(Mz_{m})^{-\beta-d}\sum_{t=T_{m-1}+1}^{T_{m}}\sum_{j\in S_{m}}\frac{1}{2^{Z}}\sum_{\omega_{-(m,j)}\in\Omega_{-(m,j)}}\underbrace{\sum_{l\in\{\pm 1\}}\mathbb{E}_{\pi,\omega_{m,j}=l}^{t}P_{X}(\pi_{t}(X_{t})\neq l\mid X_{t}\in B_{m,j})}_{U_{m,j}^{t}}.$ (34)
Here, step (i) uses the fact that regret is only incurred on $B_{m,j}$’s and the optimal action is specified by $\omega_{m,j}$; we use $\omega_{-(m,j)}$ to represent the vector after leaving out the $j$-th entry in the $m$-th block of $\omega$. By Le Cam’s method, one has
$U_{m,j}^{t}$ $\geq 1-\|\mathbb{P}_{\pi,\omega_{m,j}=-1}^{t}-\mathbb{P}_{\pi,\omega_{m,j}=1}^{t}\|_{\mathrm{TV}}$
$\geq 1-\|\mathbb{P}_{\pi,\omega_{m,j}=-1}^{T_{m}}-\mathbb{P}_{\pi,\omega_{m,j}=1}^{T_{m}}\|_{\mathrm{TV}}$
$=\int\min\left\{\mathrm{d}\mathbb{P}_{\pi,\omega_{m,j}=-1}^{T_{m}},\mathrm{d}\mathbb{P}_{\pi,\omega_{m,j}=1}^{T_{m}}\right\}$
$\geq\int_{A_{m}}\min\left\{\mathrm{d}\mathbb{P}_{\pi,\omega_{m,j}=-1}^{T_{m}},\mathrm{d}\mathbb{P}_{\pi,\omega_{m,j}=1}^{T_{m}}\right\},$
where the second inequality is due to $t\leq T_{m}$. Since the available observations for $\pi$ at $T_{m}$ are the same as those at $T_{m-1}$ under $A_{i}$, we continue to lower bound
$U_{m,i,j}^{t}$ $\geq\int_{A_{m}}\min\left\{\mathrm{d}\mathbb{P}_{\pi,\omega_{m,j}=-1}^{T_{m-1}},\mathrm{d}\mathbb{P}_{\pi,\omega_{m,j}=1}^{T_{m-1}}\right\}$
$=\frac{1}{2}\int_{A_{m}}\left(\mathrm{d}\mathbb{P}_{\pi,\omega_{m,j}=-1}^{T_{m-1}}+\mathrm{d}\mathbb{P}_{\pi,\omega_{m,j}=1}^{T_{m-1}}-|\mathrm{d}\mathbb{P}_{\pi,\omega_{m,j}=-1}^{T_{m-1}}-\mathrm{d}\mathbb{P}_{\pi,\omega_{m,j}=1}^{T_{m-1}}|\right)$
$\geq\frac{1}{2}\left(\mathbb{P}_{\pi,\omega_{m,j}=-1}^{T_{m-1}}(A_{m})+\mathbb{P}_{\pi,\omega_{m,j}=1}^{T_{m-1}}(A_{m})\right)-\|\mathbb{P}_{\pi,\omega_{m,j}=-1}^{T_{m-1}}-\mathbb{P}_{\pi,\omega_{m,j}=1}^{T_{m-1}}\|_{\mathrm{TV}}$
$\geq\frac{1}{2}\left(\mathbb{P}_{\pi,\omega_{m,j}=-1}(A_{m})+\mathbb{P}_{\pi,\omega_{m,j}=1}(A_{m})\right)-\frac{1}{2M},$
where the last step applies Lemma 2.
Plugging the above back to (34), we obtain
$\sup_{f\in\mathcal{F}}$ $R_{T}(\pi,f)$
$\geq D_{\phi}(Mz_{m})^{-(\beta+d)}\sum_{t=T_{m-1}+1}^{T_{m}}\sum_{j\in S_{m}}\frac{1}{2^{Z+1}}\sum_{\omega_{-(m,j)}\in\Omega_{-(m,j)}}\left(\mathbb{P}_{\pi,\omega_{m,j}=-1}(A_{m})+\mathbb{P}_{\pi,\omega_{m,j}=1}(A_{m})-\frac{1}{M}\right)$
$=D_{\phi}(Mz_{m})^{-(\beta+d)}\sum_{t=T_{m-1}+1}^{T_{m}}\sum_{j\in S_{m}}\frac{1}{2}\left(\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{P}_{\pi,\omega}(A_{m})-\frac{1}{2M}\right)$
$=\frac{1}{2}D_{\phi}(Mz_{m})^{-(\beta+d)}(T_{m}-T_{m-1})s_{m}\left(\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{P}_{\pi,\omega}(A_{m})-\frac{1}{2M}\right).$
Since $\sum_{k=1}^{M}\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{P}_{\pi,\omega}(A_{k})\geq 1$, there exists some $m^{\star}\in[M]$ such that $\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{P}_{\pi,\omega}(A_{m^{\star}})\geq 1/M$. When $m^{\star}=1$, one has,
$\sup_{f\in\mathcal{F}}R_{T}(\pi,f)\apprge M^{-3}\cdot T_{1}\asymp M^{-3}\cdot T^{h_{M}(\alpha)}$
When $m^{\star}\geq 2$,
$\sup_{f\in\mathcal{F}}R_{T}(\pi,f)$ $\apprge M^{-2}T_{m^{\star}}(Mz_{m^{\star}})^{-\beta(1+\alpha)}$
$\asymp M^{-2}T_{m^{\star}}(M^{2}T_{m^{\star}-1})^{-\gamma(\alpha)}\apprge M^{-4}\cdot T^{h_{M}(\alpha)}.$
Appendix B Detailed analysis of the variational problem
We recall that
$\psi_{M}(\bm{u})\;=\;\sup_{\alpha\in\mathcal{K}}\Psi_{M}(u,\alpha),$
where
$\Psi_{M}(\bm{u},\alpha)=\max\Bigl\{\,u_{1},\ u_{2}-\gamma(\alpha)u_{1},\ \dots,\ u_{M-1}-\gamma(\alpha)u_{M-2},\ 1-\gamma(\alpha)u_{M-1}\Bigr\}-h_{M}(\alpha),$
with $\gamma(\alpha)=\tfrac{(\alpha+1)\beta}{2\beta+d}$ for $\alpha\in[0,d/\beta]$, and $\Psi_{M}(\bm{u},\infty)=u_{1}$. Here,
$\bm{u}\in\mathcal{U}_{M}=\{\bm{u}\in\mathbb{R}^{M-1}:0\leq u_{1}\leq\cdots\leq u_{M-1}\leq 1\}.$
In this section, we collect several useful facts of the variational problem.
Convexity of $\psi_{M}(\bm{u})$.
For each fixed $\alpha\in\mathcal{K}$, the payoff function $\Psi_{M}(\bm{u},\alpha)$ is piecewise linear, and hence convex in $\bm{u}$. As a result, $\psi_{M}(\bm{u})=\sup_{\alpha\in\mathcal{K}}\Psi_{M}(\bm{u},\alpha)$ is a convex function.
Existence of minimizer.
Note that $\Psi_{M}(\bm{u},\alpha):\mathcal{U}_{M}\times\mathcal{K}\rightarrow\mathbb{R}$ is jointly continuous in $\bm{u}$ and $\alpha$. We can apply Berge’s maximum theorem to show $\psi_{M}(\bm{u})=\sup_{\alpha\in\mathcal{K}}\Psi_{M}(\bm{u},\alpha)$ is continuous on $\mathcal{U}_{M}$. Consequently, by the Weierstrass extreme value theorem, there exists some $\bm{u}^{\star}\in\mathcal{U}_{M}$ such that $\psi_{M}(\bm{u}^{\star})=\psi_{M,\mathcal{K}}^{\star}$.
Positive optimal value.
We know that for every $\alpha\in\mathcal{K}$,
$\inf_{\bm{u}\in\mathcal{U}_{M}}\Psi_{M}(\bm{u},\alpha)=0,$
and 0 is achievable by some $\bm{u}^{\star}(\alpha)\in\mathcal{U}_{M}$. We also know that for any $\alpha_{1}\neq\alpha_{2}\in\mathcal{K}$, $\bm{u}^{\star}(\alpha_{1})\neq\bm{u}^{\star}(\alpha_{2})$.
Now suppose that $\psi_{M}^{\star}=0$, and let $\bm{u}^{\star}$ be the minimizer, whose existence has been shown above. Then we have
$\psi_{M}(\bm{u}^{\star})=\sup_{\alpha\in\mathcal{K}}\Psi_{M}(\bm{u}^{\star},\alpha)=0.$
That is, for every $\alpha\in\mathcal{K}$, we have $\Psi_{M}(\bm{u}^{\star},\alpha)\leq 0$. Taking the previous displays together, we arrive at the conclusion that
$\Psi_{M}(\bm{u}^{\star},\alpha)=0,\qquad\text{for all }\alpha\in\mathcal{K}.$
However, this contradicts with the fact that for different $\alpha$’s, we have different minimizers. As a result, we necessarily have $\psi_{M}^{\star}>0$.
Subdifferential.
By the rule of the subdifferential, we know that
$\partial\psi_{M}(\bm{u})=\operatorname{conv}\Biggl(\;\bigcup_{\alpha\in\mathcal{A}(\bm{u})}\ \partial_{\bm{u}}\Psi_{M}(\bm{u},\alpha)\;\Biggr),$
where $\mathcal{A}(\bm{u})=\{\alpha:\Psi_{M}(\bm{u},\alpha)=\psi_{M}(\bm{u})\}$ denotes the set of active maximizers in the $\sup$.
Now we move on to $\partial_{\bm{u}}\Psi_{M}(\bm{u},\alpha)$. For each $\alpha<\infty$ the inner maximum has affine pieces with gradients
$\bm{g}_{1}=\bm{e}_{1},\qquad\bm{g}_{i}(\gamma)=\bm{e}_{i}-\gamma\bm{e}_{i-1}\quad(i=2,\dots,M-1),\qquad\bm{g}_{M}(\gamma)=-\gamma\bm{e}_{M-1},$
where $\bm{e}_{i}$ is the $i$-th standard basis vector in $\mathbb{R}^{M-1}$. At $\alpha=\infty$ only the block $u_{1}$ is active, with gradient $\bm{g}_{1}=\bm{e}_{1}$. Therefore,
$\partial\psi_{M}(\bm{u})=\operatorname{conv}\Bigl\{\,\bm{g}_{i}(\gamma(\alpha)):\ \alpha\in\mathcal{A}(\bm{u}),\ i\in\mathcal{I}(\bm{u},\alpha)\,\Bigr\},$
where $\mathcal{I}(\bm{u},\alpha)$ is the set of indices $i$ attaining the max in $\Psi_{M}(\bm{u},\alpha)$.
Carathéodory’s theorem in $\mathbb{R}^{M-1}$ implies that any point of $\partial\psi_{M}(\bm{u})$ can be represented as a convex combination of at most $M$ vectors. Concretely, for any $\bm{v}\in\partial\psi_{M}(\bm{u})$ there exist pairs $(\alpha_{k},i_{k})$ with $\alpha_{k}\in\mathcal{A}(\bm{u})$, $i_{k}\in\mathcal{I}(\bm{u},\alpha_{k})$ and weights $\theta_{k}\geq 0$, $\sum_{k=1}^{M}\theta_{k}=1$, such that
$\bm{v}=\sum_{k=1}^{M}\theta_{k}\,\bm{g}_{i_{k}}(\gamma(\alpha_{k})).$ (35)
Note that if $\alpha_{k}=\infty$, we must have $i_{k}=1$.
We record a useful property of this subdifferential.
Lemma 9.
For any $\bm{u}$ with $\mathbf{0}\in\partial\psi_{M}(\bm{u})$, we have for each $1\leq i\leq M$, there exists some $\alpha_{i}\in\mathcal{K}$ such that $\eta_{i}(\bm{u},\alpha_{i})-h_{M}(\alpha_{i})=\psi_{M}(\bm{u})$.
Proof.
By equation (35), there exist pairs $(\alpha_{k},i_{k})$ with $\alpha_{k}\in\mathcal{A}(\bm{u})$, $i_{k}\in\mathcal{I}(\bm{u},\alpha_{k})$ and weights $\theta_{k}\geq 0$, $\sum_{k=1}^{M}\theta_{k}=1$, such that
$\bm{0}=\sum_{k=1}^{M}\theta_{k}\,\bm{g}_{i_{k}}(\gamma(\alpha_{k})).$
Since $\sum_{k=1}^{M}\theta_{k}=1$, there exists some $1\leq k\leq M$ such that $\theta_{k}>0$. Let $i_{k}$ for the corresponding index for $\bm{g}$, i.e., $\bm{g}_{i_{k}}$ is included in the convex combination. Suppose that $i_{k}=1$. By the structure of $\bm{g}_{1}$, we know that $\theta_{k}\bm{g}_{1}$ is positive in the first entry. To cancel this, we must have $\bm{g}_{2}$ in the convex combination, which further brings $\bm{g}_{3}$ into the convex combination. Chaining this argument, we arrive at the conclusion that all $\{\bm{g}_{i}\}_{1\leq i\leq M}$ must be involved in the convex combination. The argument continues to hold if $i_{k}\geq 2$.
Since the set $\{i_{k}\}_{1\leq k\leq M}=\{1,2,\ldots,M\}$, by the definition of $(\alpha_{k},i_{k})$, we know that for each $1\leq i\leq M$, we have some $\alpha_{i}\in\mathcal{K}$ such that $\eta_{i}(\bm{u},\alpha_{i})-h_{M}(\alpha_{i})=\psi_{M}(\bm{u})$. ∎
Explicit conic representation of $N_{\mathcal{U}}(\bm{u})$.
Define $u_{0}\coloneqq 0$ and $u_{M}\coloneqq 1$. For $i=1,\dots,M$, set
$\bm{d}_{i}\;\coloneqq\;\bm{e}_{i-1}-\bm{e}_{i}\in\mathbb{R}^{M-1},\qquad\text{with the convention }\bm{e}_{0}\coloneqq\bm{0},\ \bm{e}_{M}\coloneqq\bm{0}.$
Then define the active set at $\bm{u}$ as
$I(\bm{u})\;\coloneqq\;\{\,i\in\{1,\dots,M\}:\ u_{i}=u_{i-1}\,\}.$
The normal cone is the conic hull of the active normals:
$N_{\mathcal{U}}(\bm{u})\;=\;\Bigl\{\,\bm{n}\in\mathbb{R}^{M-1}:\ \bm{n}=\sum_{i\in I(\bm{u})}\lambda_{i}\,\bm{d}_{i},\ \ \lambda_{i}\geq 0\Bigr\}.$ (36)
The first-order optimality condition.
The KKT condition  $\bm{0}\in\partial\psi_{M}(\bm{u}^{\star})+N_{\mathcal{U}}(\bm{u}^{\star})\,$ is therefore equivalent to the existence of multipliers $\{\lambda_{i}\}_{i=1}^{M}$ with $\lambda_{i}\geq 0$ and $\lambda_{i}=0$ if $i\notin I(\bm{u}^{\star})$, and weights $\{\theta_{k}\}_{k=1}^{M}$ as above, such that
$\underbrace{\sum_{k=1}^{M}\theta_{k}\,\bm{g}_{i_{k}}\bigl(\gamma(\alpha_{k})\bigr)}_{\eqqcolon\bm{v}}\;+\;\underbrace{\sum_{i\in I(\bm{u}^{\star})}\lambda_{i}\,(\bm{e}_{i-1}-\bm{e}_{i})}_{\eqqcolon\bm{n}}\;=\;\bm{0}.$
Now, we are ready to establish an important property about the variational problem.
B.1 The minimizer lies in the interior
[…]
For the sake of contradiction, assume that $\bm{u}^{\star}$ is a minimizer lying on the boundary of $\mathcal{U}_{M}$. By definition, there exists some index $1\leq j\leq m$ such that $u^{\star}_{j}=u^{\star}_{j-1}$. Here we again implicitly define $u^{\star}_{0}=0$, and $u^{\star}_{M}=1$. Consequently, we have the following lemma.
Lemma 10.
Let $\bm{u}^{\star}$ be a minimizer. Suppose that $u^{\star}_{j}=u^{\star}_{j-1}$ for some $1\leq j\leq M$, then for any $\alpha\in\mathcal{K}$, we have the inequality
$\eta_{j}(\bm{u}^{\star},\alpha)-h_{M}(\alpha)<\psi_{M}(\bm{u}^{\star}).$
Proof.
We consider the following two cases.
Case 1: $u^{\star}_{j-1}=0$.
In this case, for any $\alpha\in\mathcal{K}$, we have $\eta_{j}(\bm{u}^{\star},\alpha)-h_{M}(\alpha)=u^{\star}_{j}-\gamma(\alpha)u^{\star}_{j-1}-h_{M}(\alpha)=-h_{M}(\alpha)\leq 0$, while $\psi_{M}(\bm{u}^{\star})=\psi_{M}^{\star}>0$. Hence the desired inequality holds.
Case 2: $u^{\star}_{j-1}>0$.
Let $k\geq 0$ be the largest index such that $\bm{u}^{\star}_{k}<\bm{u}^{\star}_{j-1}$. Such $k$ is guaranteed to exist because $\bm{u}^{\star}_{0}=0<\bm{u}^{\star}_{j-1}$. In this case, we have In other words, $\bm{u}^{\star}_{k}=\bm{u}^{\star}_{i-1}$ for $j+1\leq k\leq i-1$. One has
$\eta_{j}(\bm{u}^{\star},\alpha)=(1-\gamma(\alpha))\bm{u}^{\star}_{j-1},$
while
$\eta_{k+1}(\bm{u}^{\star},\alpha)=\bm{u}^{\star}_{k+1}-\gamma(\alpha)\bm{u}^{\star}_{k}=\bm{u}^{\star}_{j-1}-\gamma(\alpha)\bm{u}^{\star}_{k}>(1-\gamma(\alpha))\bm{u}^{\star}_{j-1}=\eta_{j}(\bm{u}^{\star},\alpha).$
Here, the inequality is due to $\bm{u}^{\star}_{k}<\bm{u}^{\star}_{i-1}$. Therefore, we have $\eta_{j}(\bm{u}^{\star},\alpha)-h_{M}(\alpha)<\eta_{k+1}(\bm{u}^{\star},\alpha)-h_{M}(\alpha)\leq\psi_{M}(\bm{u}^{\star})$. ∎
Combining Lemma 10 and Lemma 9, we see that $\mathbf{0}\notin\partial\psi_{M}(\bm{u}^{\star})$, $\bm{v}\neq\bm{0}$. As a result, $\bm{n}\neq\bm{0}$. Let $j$ be the smallest index in $I(\bm{u}^{\star})$ such that $\lambda_{j}>0$. By Lemma 10 again, we know that $i_{k}\neq j$ for all $k$’s in the convex combination representation of $\bm{v}$.
First, suppose that $j\geq 2$. Consider the coordinate $v_{j-1}+n_{j-1}$. By the definition of the normal vector, we know that $n_{j-1}>0$. Since $i_{k}\neq j$ for all $k$’s, we also know that $v_{j-1}\geq 0$. This contradicts with $\bm{v}+\bm{n}=\bm{0}$.
Second, suppose that $j=1$, i.e., $u^{\star}_{1}=0$, and $\lambda_{1}>0$. If $\lambda_{2}=0$, then we must have $n_{1}<0$. Note that $i_{k}\neq 1$ for all $k$, and hence $v_{1}\leq 0$. This contradicts with the first-order optimality condition. Consequently, we must have $\lambda_{2}>0$. Now consider the second coordinate $v_{2}+n_{2}$. If $\lambda_{3}=0$, then $n_{2}<0$. However $i_{k}\neq 2$ for all $k$, and hence $v_{2}\leq 0$. As a result, we can only have $\lambda_{3}>0$. Continuing this argument, we must have $\lambda_{j}>0$ for all $1\leq j\leq M$. In other words, $u^{\star}_{j}=u^{\star}_{j-1}$ for all $1\leq j\leq M$, which is impossible.
In all, we have proved via contradiction that $\bm{u}^{\star}$ must lie in the interior of the feasible set $\mathcal{U}_{M}$.
B.2 Reduction to a system of equations
The analysis carried out so far paves the way for studying the original problem in an alternative form, which proves much more convenient for later use.
Since $\bm{u}^{\star}$ is a minimizer lying in the interior of $\mathcal{U}_{M}$, we have $\mathbf{0}\;\in\;\partial\psi_{M}(\bm{u}^{\star})$. By Lemma 9, for each $1\leq i\leq M$, there exists some $\alpha_{i}\in\mathcal{K}$ such that $\eta_{i}(\bm{u}^{\star},\alpha_{i})-h_{M}(\alpha_{i})=\psi_{M}(\bm{u}^{\star})$. For $i\geq 2$, we have
$\eta_{i}(\bm{u}^{\star},\alpha_{i})-h_{M}(\alpha_{i})=u^{\star}_{i}-\gamma(\alpha)u^{\star}_{i-1}-h_{M}(\alpha_{i})=\psi_{M}(\bm{u}^{\star})>0.$
It is clear that $\alpha_{i}<\infty$. Otherwise the equality would not hold.
Denote by $\mathcal{S}=[\beta/(2\beta+d),(\beta+d)/(2\beta+d)]=[\gamma_{\mathrm{min}},\gamma_{\mathrm{max}}]$ the feasible range of $\gamma(\alpha)$. From now on, we redefine $h_{M}:\mathcal{S}\rightarrow\mathbb{R}$ as
$h_{M}(\gamma)=\frac{1-\gamma}{1-\gamma^{M}}.$
To avoid notation cluster, we write $\bm{u}$ for $\bm{u}^{\star}$. By the optimality condition,
$u_{1}$ $=\max_{\gamma\in\mathcal{S}}u_{i}-\gamma u_{i-1}-h_{M}(\gamma),\quad 2\leq i\leq M-1$
$=\max_{\gamma\in\mathcal{S}}1-\gamma u_{M-1}-h_{M}(\gamma).$ (37)
Define the function $\phi_{M}(x)=\min_{\gamma\in\mathcal{S}}\gamma x+h_{M}(\gamma)$ for $x\in(0,1)$. Rearranging the above equations, we have
$u_{i}=u_{1}+\phi_{M}(u_{i-1}),\quad 2\leq i\leq M,$ (38)
where we write $u_{M}=1$ for convenience. Denote by $\gamma_{i}=\operatorname*{arg\,min}_{\gamma\in\mathcal{S}}\gamma u_{i-1}+h_{M}(\gamma)$. Clearly, one has $\gamma_{i}=\beta(1+\alpha_{i})/(2\beta+d)$.
The system of equations in (B.2) is an important consequence of the optimality condition. From now on, we will focus on this system rather than the original objective function. As we shall soon see, it allows us to establish several interesting properties about the minimizer $\bm{u}$ and the sequence $\{\gamma_{i}\}$.
B.2.1 Monotonicity of $\{\gamma_{i}\}$
First, we show the sequence $\{\gamma_{i}\}$ is non-increasing, which in turn translates to the monotonicity of $\alpha_{i}$. Let $\eta_{v}(\gamma)=v\cdot\gamma+h_{M}(\gamma)$. Since $h_{M}^{\prime}(\gamma)<0$ and $h_{M}^{\prime\prime}(\gamma)>0$,
$\eta^{\prime}(\gamma)=v+h_{M}^{\prime}(\gamma)$
is strictly increasing in $\gamma$, so $\eta$ is strictly convex and has a unique minimizer $\gamma^{*}(v)$.
Define the thresholds
$v_{L}\coloneqq-h_{M}^{\prime}(\gamma_{\mathrm{max}}),\quad v_{U}\coloneqq-h_{M}^{\prime}(\gamma_{\mathrm{min}})\quad\text{with }0<v_{L}<v_{U}.$
Then
$\gamma^{*}(v)=\begin{cases}\gamma_{\mathrm{max}},&0<v<c_{L},\\ \textrm{the solution to }-h_{M}^{\prime}(\gamma)=v,&v_{L}\leq v\leq v_{U},\\ \gamma_{\mathrm{min}},&v>v_{U}.\end{cases}$
Because $h_{M}^{\prime}(\gamma)$ is decreasing in $\gamma$, the solution of $h_{M}^{\prime}(\gamma)=v$ becomes smaller when $v$ increases. Thus $\gamma^{*}(v)$ is nonincreasing in $v$: it is at $\gamma_{\mathrm{max}}$ for small $v$, moves left continuously through the interior as $v$ grows, and sticks at $\gamma_{\mathrm{min}}$ for large $v$. Since $\{u_{i}\}_{i=1}^{M-1}$ is an increasing sequence, the sequence $\{\gamma^{*}(u_{i})\}_{i=1}^{M}$ is non-increasing.
B.2.2 Behavior of the individual $u_{i}$
The first lemma provides lower bound to $u_{1}$. For $a\in(0,1)$ and $n\geq 2$, define $S_{n}(a)\coloneqq\sum_{k=0}^{n-1}a^{-k}$.
Lemma 11.
Fix any $c\in\mathcal{S}$. The first component $u_{1}$ is lower bounded by
$u_{1}\geq\frac{c^{-1}\gamma_{\mathrm{max}}{}^{-(M-3)}\bigl(1-h_{M}(c)\bigr)-h_{M}(\gamma_{\mathrm{max}})S_{M-2}(\gamma_{\mathrm{max}})}{S_{M-2}(\gamma_{\mathrm{max}})+\gamma_{\mathrm{max}}+c^{-1}\gamma_{\mathrm{max}}{}^{-(M-3)}}.$
Proof.
In view of the optimality condition (B.2), we have
$u_{1}$ $\geq u_{j+1}-\gamma_{\mathrm{max}}u_{j}-h_{M}(\gamma_{\mathrm{max}}),\quad 1\leq j\leq M-2,$
$u_{1}$ $\geq 1-cu_{M-1}-h_{M}(c).$ (39)
Multiplying the $j$-th inequality by $\gamma_{\mathrm{max}}^{-(j-1)}$ and summing over $1\leq j\leq M-2$, we obtain
$u_{1}\sum_{j=1}^{M-2}\gamma_{\mathrm{max}}^{-(j-1)}\geq\gamma_{\mathrm{max}}^{-(M-3)}u_{M-1}-\gamma_{\mathrm{max}}u_{1}-h_{M}(\gamma_{\mathrm{max}})\sum_{j=1}^{M-2}\gamma_{\mathrm{max}}^{-(j-1)}.$
Recall that $S_{n}(\gamma_{\mathrm{max}})=\sum_{k=0}^{n-1}\gamma_{\mathrm{max}}^{-k}$. We have $\sum_{j=1}^{M-2}\gamma_{\mathrm{max}}^{-(j-1)}=\sum_{j=0}^{M-3}\gamma_{\mathrm{max}}^{-j}=S_{M-2}(\gamma_{\mathrm{max}}).$ Use this to rewrite the inequality as
$u_{1}(S_{M-2}(\gamma_{\mathrm{max}})+\gamma_{\mathrm{max}})$ $\geq\gamma_{\mathrm{max}}^{-(M-3)}u_{M-1}-h_{M}(\gamma_{\mathrm{max}})S_{M-2}(\gamma_{\mathrm{max}})$
$\geq\gamma_{\mathrm{max}}^{-(M-3)}c^{-1}(1-u_{1}-h_{M}(c))-h_{M}(\gamma_{\mathrm{max}})S_{M-2}(\gamma_{\mathrm{max}})$
$=-c^{-1}\gamma_{\mathrm{max}}^{-(M-3)}u_{1}+c^{-1}\gamma_{\mathrm{max}}^{-(M-3)}(1-h_{M}(c))-h_{M}(\gamma_{\mathrm{max}})S_{M-2}(\gamma_{\mathrm{max}}),$
where the second step is due to relation (39). Combining terms we reach
$u_{1}(S_{M-2}(\gamma_{\mathrm{max}})+\gamma_{\mathrm{max}}+c^{-1}\gamma_{\mathrm{max}}^{-(M-3)})\geq c^{-1}\gamma_{\mathrm{max}}^{-(M-3)}(1-h_{M}(c))-h_{M}(\gamma_{\mathrm{max}})S_{M-2}(\gamma_{\mathrm{max}}).$
Rearranging terms yields the desired claim. ∎
Lemma 11 lower bounds the value of the first component $u_{1}$. Since $u_{1}=\psi_{M}^{\star}$ by the optimality condition, it provides a lower bound for the optimal objective value as well.
Similarly, we have the upper bound on $u_{1}$ in the following lemma.
Lemma 12.
We have
$u_{1}\leq\frac{(M+1)\gamma_{\max}^{M-1}}{(1-\gamma_{\max})^{2}}.$
Proof.
The key identity to establish an upper bound on $u_{1}$ is
$u_{1}+\phi_{M}(u_{M-1})=u_{1}+\inf_{\gamma\in\mathcal{S}}u_{M-1}\gamma+h_{M}(\gamma)=1.$
Now we split the proof into two cases: (1) when $\gamma_{M}=\gamma_{\max}$, and (2) when $\gamma_{M}<\gamma_{\max}$.
Case 1: when $\gamma_{M}=\gamma_{\max}$.
In this case, using the fact that $\gamma_{i}$ is monotonically decreasing, we know that
$\gamma_{2}=\gamma_{3}=\cdots=\gamma_{M}=\gamma_{\max}.$
This actually allows us to solve for $u_{1}$ exactly:
$u_{1}=\frac{\gamma_{\max}^{M-1}(1-\gamma_{\max})^{2}}{(1-\gamma_{\max}^{M})^{2}}.$
Case 2: when $\gamma_{M}<\gamma_{\max}$.
In this case, we have the relationship
$u_{M-1}\geq-h_{M}^{\prime}(\gamma_{M}).$
This together with the key identity yields
$u_{1}$ $=1-\inf_{\gamma\in\mathcal{S}}u_{M-1}\gamma+h_{M}(\gamma)$
$=1-u_{M-1}\gamma_{M}-h_{M}(\gamma_{M})$
$\leq 1+h_{M}^{\prime}(\gamma_{M})\gamma_{M}-h_{M}(\gamma_{M})$
$\leq 1+h_{M}^{\prime}(\gamma_{\max})\gamma_{\max}-h_{M}(\gamma_{\max}),$
where the last steps uses the fact that $h_{M}^{\prime}(\gamma)\gamma-h_{M}(\gamma)$ is increasing in $\gamma$. Write this upper bound explicitly to see that
$u_{1}\leq\frac{\gamma_{\max}^{M}(M(1-\gamma_{\max})-1+\gamma_{\max}^{M})}{(1-\gamma_{\max}^{M})^{2}}.$
In both cases, the upper bound can be further relaxed to the one stated in the lemma. ∎
The next two lemmas are about the gaps between consecutive $u_{i-1}$ and $u_{i}$.
Lemma 13.
For any $M\geq 3$, one has
$u_{i-1}\leq u_{i}\cdot(\gamma_{i+1}+\frac{1}{2}),\quad\forall 2\leq i\leq M-1.$
See Appendix B.3 for the proof of Lemma 13.
Denote by $\Delta_{i}=u_{i+1}-u_{i}$.
Lemma 14.
One has $\gamma_{M}\leq\Delta_{M-1}/\Delta_{M-2}\leq\gamma_{M-1}$.
Proof.
By concavity of $\phi_{M}(\cdot)$,
$\phi_{M}^{\prime}(u_{M-1})\leq\frac{\phi_{M}(u_{M-1})-\phi_{M}(u_{M-2})}{u_{M-1}-u_{M-2}}\leq\phi_{M}^{\prime}(u_{M-2}).$
By relation (38), we have $\phi_{M}(u_{M-1})-\phi_{M}(u_{M-2})=u_{M}-u_{M-1}=\Delta_{M-1}$. Meanwhile, $\phi_{M}^{\prime}(u_{M-1})=\gamma_{M}$ and $\phi_{M}^{\prime}(u_{M-2})=\gamma_{M-1}$. Hence, $\gamma_{M}\leq\Delta_{M-1}/\Delta_{M-2}\leq\gamma_{M-1}$. ∎
B.2.3 Monotonicity of $\psi_{M}^{\star}$
We will show $\psi_{M}^{\star}$ is strictly decreasing in $M$ when $M$ is small, which proves Proposition 7.
Let $\bm{u}^{(M)}=\operatorname*{arg\,min}_{\bm{v}\in\mathcal{U}_{M}}\sup_{\alpha\in\mathcal{K}}\Psi_{M}(\bm{v},\alpha)$ and $\bm{u}^{(M+1)}=\operatorname*{arg\,min}_{\bm{v}\in\mathcal{U}_{M+1}}\sup_{\alpha\in\mathcal{K}}\Psi_{M+1}(\bm{v},\alpha)$. By the optimality condition, we have $\psi_{M}^{\star}=u_{1}^{(M)}$ and $\psi_{M+1}^{\star}=u_{1}^{(M+1)}$. It is equivalent to show $u_{1}^{(M+1)}<u_{1}^{(M)}$.
We first give a high-level description of the proof idea. For the sake of contradiction, suppose that the optimal solution for $M+1$ has $u_{1}^{(M+1)}\geq u_{1}^{(M)}$. We can then iteratively solve the recursion in (38) and obtain $u_{2}^{(M+1)},\ldots,u_{M}^{(M+1)}$. Then we show that it must contradict the equation
$1=u_{1}^{(M+1)}+\phi_{M+1}(u_{M}^{(M+1)}).$
By the definition of $\phi_{M+1}$, it suffices to prove that for all $\gamma\in S$, we have
$u_{1}^{(M+1)}+\gamma u_{M}^{(M+1)}+h_{M+1}(\gamma)-1>0.$ (40)
We therefore need a lower bound on $u_{M}^{(M+1)}$ and also a lower bound on $u_{1}^{(M+1)}$. To lower bound $u_{1}^{(M)}$, we can apply Lemma 11 with $c=\gamma_{\mathrm{max}}$,
$u_{1}^{(M)}\geq\frac{\gamma_{\mathrm{max}}^{M-1}}{S_{M}^{2}(\gamma_{\mathrm{max}})}.$
Next, we turn to lower bounding $u_{M}^{(M+1)}$. By the recursive relation in (38), we know that
$u_{k+1}^{(M+1)}\;=\;u_{1}^{(M+1)}\;+\;\phi_{M+1}\!\big(u_{k}^{(M+1)}\big).$
Denote by $\Delta_{M}(\gamma)=h_{M}(\gamma)-h_{M+1}(\gamma)$. For $k=1$, one has
$u_{2}^{(M+1)}$ $\;=\;u_{1}^{(M+1)}\;+\;\phi_{M+1}\!\big(u_{1}^{(M+1)}\big)$
$\;=\;u_{1}^{(M+1)}\;+\;\inf_{\gamma}\!\big(\gamma u_{1}^{(M+1)}+h_{M+1}(\gamma)\big)$
$\geq\;u_{1}^{(M)}\;+\;\inf_{\gamma}\!\big(\gamma u_{1}^{(M)}+h_{M}(\gamma)+h_{M+1}(\gamma)-h_{M}(\gamma)\big)$
$\geq u_{1}^{(M)}+\inf_{\gamma}(\gamma u_{1}^{(M)}+h_{M}(\gamma))-\Delta(\gamma_{\mathrm{max}})$
$=u_{2}^{(M)}-\Delta(\gamma_{\mathrm{max}}),$
where the first inequality is due to the assumption $u_{1}^{(M+1)}\geq u_{1}^{(M)}$, and the second inequality is because the function $\Delta_{M}(\cdot)$ is increasing. Similarly when $k=2$, we have
$u_{3}^{(M+1)}$ $\;=\;u_{1}^{(M+1)}\;+\;\phi_{M+1}\!\big(u_{2}^{(M+1)}\big)$
$\;=\;u_{1}^{(M+1)}\;+\;\inf_{\gamma}\!\big(\gamma u_{2}^{(M+1)}+h_{M+1}(\gamma)\big)$
$=\;u_{1}^{(M+1)}\;+\;\inf_{\gamma}\!\big(\gamma u_{2}^{(M)}-\gamma u_{2}^{(M)}+\gamma u_{2}^{(M+1)}+h_{M}(\gamma)+h_{M+1}(\gamma)-h_{M}(\gamma)\big)$
$\geq u_{1}^{(M)}+\inf_{\gamma}(\gamma u_{2}^{(M)}+h_{M}(\gamma))-\Delta(\gamma_{\mathrm{max}})-\gamma_{\mathrm{max}}(u_{2}^{(M)}-u_{2}^{(M+1)})$
$=u_{3}^{(M)}-(1+\gamma_{\mathrm{max}})\Delta(\gamma_{\mathrm{max}}).$
Recursively, we obtain
$u_{M}^{(M+1)}\geq u_{M}^{(M)}-S_{M-2}(\gamma_{\mathrm{max}})\Delta(\gamma_{\mathrm{max}})=1-S_{M-2}(\gamma_{\mathrm{max}})\Delta(\gamma_{\mathrm{max}}).$
Substituting back into the key equation we aim to prove to see that
$u_{1}^{(M+1)}+\gamma u_{M}^{(M+1)}+h_{M+1}(\gamma)-1$ $\geq u_{1}^{(M)}+\gamma(1-S_{M-2}(\gamma_{\mathrm{max}})\Delta(\gamma_{\mathrm{max}}))+h_{M+1}(\gamma)-1$
$\geq\frac{\gamma_{\mathrm{max}}^{M-1}}{S_{M}^{2}(\gamma_{\mathrm{max}})}+\gamma(1-S_{M-2}(\gamma_{\mathrm{max}})\Delta(\gamma_{\mathrm{max}}))+h_{M+1}(\gamma)-1.$
For $M=2,3,4$, the RHS of the above relation simplifies to
$M=2:\quad$ $\frac{\gamma_{\mathrm{max}}}{(1+\gamma_{\mathrm{max}})^{2}}+\gamma+\frac{1}{1+\gamma+\gamma^{2}}-1=\frac{\gamma_{\mathrm{max}}}{(1+\gamma_{\mathrm{max}})^{2}}+\frac{\gamma^{3}}{1+\gamma+\gamma^{2}}>0,$
$M=3:\quad$ $\frac{\gamma_{\mathrm{max}}^{2}}{(1+\gamma_{\mathrm{max}}+\gamma_{\mathrm{max}}^{2})^{2}}+\gamma\Big(1-\frac{\gamma_{\mathrm{max}}^{3}}{S_{3}(\gamma_{\mathrm{max}})S_{4}(\gamma_{\mathrm{max}})}\Big)+\frac{1}{S_{4}(\gamma)}-1>0,$
$M=4:\quad$ $\frac{\gamma_{\mathrm{max}}^{3}}{S_{4}(\gamma_{\mathrm{max}})^{2}}+\gamma\Big(1-\frac{S_{2}(\gamma_{\mathrm{max}})\,\gamma_{\mathrm{max}}^{4}}{S_{4}(\gamma_{\mathrm{max}})S_{5}(\gamma_{\mathrm{max}})}\Big)+\frac{1}{S_{5}(\gamma)}-1>0,$
for all $\gamma\in\mathcal{S}$. Consequently, relation (40) holds and we reach $u_{1}^{(M+1)}+\phi_{M+1}(u_{M}^{(M+1)})>1$, which is a contradiction. Therefore, one has $u_{1}^{(M+1)}<u_{1}^{(M)}$ when $M=2,3,4$.
B.3 Proof of Lemma 13
On a high level, the proof first establishes $u_{i-1}\leq u_{i}\cdot(\gamma_{i+1}+\frac{1}{2})$ for any $2\leq i\leq M-2$ by showing $\gamma_{i+1}\geq 1/2$. To do so, it suffices to prove $\gamma_{M-1}\geq 1/2$ and use the fact that $\{\gamma_{i}\}$ is non-increasing. Next, we show that the last inequality $u_{M-2}\leq u_{M-1}\cdot(\gamma_{M}+\frac{1}{2})$ holds as well. This is achieved through analyzing different ranges of $M$ and establishing tight lower bound on $\gamma_{M}$ under that range.
B.3.1 Key lemmas
In this section, we collect several useful bounds on $u_{1}$ and $u_{M-1}$. When $\gamma_{M}$ is small, we have the following lower bound.
Lemma 15.
Fix any $c\in(\gamma_{\mathrm{min}},1/2]$. If $\gamma_{M}<c$, then
$u_{1}+\phi_{M}(u_{M-1})>\frac{c^{-1}\gamma_{\mathrm{max}}{}^{-(M-3)}\bigl(1-h_{M}(c)\bigr)-h_{M}(\gamma_{\mathrm{max}})S_{M-2}(\gamma_{\mathrm{max}})}{S_{M-2}(\gamma_{\mathrm{max}})+\gamma_{\mathrm{max}}+c^{-1}\gamma_{\mathrm{max}}{}^{-(M-3)},}-ch_{M}^{\prime}(c)+h_{M}(c).$
When $\gamma_{M-1}$ is small, we have the following lower bound.
Lemma 16.
If $\gamma_{M-1}<1/2$, then
$u_{1}+\phi_{M}(u_{M-1})>(1+\gamma_{\mathrm{min}})u_{1}+(h_{M}(\frac{1}{2})-\frac{1}{2}h_{M}^{\prime}(\frac{1}{2}))\gamma_{\mathrm{min}}+h_{M}(\gamma_{\mathrm{min}}).$
Figure 5: Lower bound of $u_{M}$ vs. batch budget $M$ when $\gamma_{M}<0.45$.
B.3.2 Main proof
Now we are ready to prove Lemma 13. We start with proving that $u_{i-1}\leq u_{i}\cdot(\gamma_{i+1}+\frac{1}{2})$ holds for all $2\leq i\leq M-2$. Then we finish with proving $u_{M-2}<u_{M-1}(\gamma_{M}+\frac{1}{2})$.
Step 1: Establishing $u_{i-1}\leq u_{i}\cdot(\gamma_{i+1}+\frac{1}{2})\;\forall 2\leq i\leq M-2$.
A key step is to prove that $\gamma_{M}\geq 0.45$ and $\gamma_{M-1}\geq 0.5$. We start by lower bounding $\gamma_{M}$ by $0.45$. Suppose that $\gamma_{M}<0.45$. Applying Lemma 15 with $c=0.45$, we have $u_{1}+\phi_{M}(u_{M-1})>1$, which contradicts with $u_{M}=-u_{1}+\phi_{M}(u_{M-1})=1$; see Figure 5. As a result, we have $\gamma_{M}\geq 0.45$
Next, we lower bound $\gamma_{M-1}$ by $0.5$. Suppose that $\gamma_{M-1}<1/2$. By Lemma 16, we have
$u_{1}+\phi_{M}(u_{M-1})>(1+\gamma_{\mathrm{min}})u_{1}+(h_{M}(\frac{1}{2})-\frac{1}{2}h_{M}^{\prime}(\frac{1}{2}))\gamma_{\mathrm{min}}+h_{M}(\gamma_{\mathrm{min}}).$
In addition, apply Lemma 11 with $c=0.45$ to obtain
$u_{1}\geq\frac{\frac{1}{0.45}\gamma_{\mathrm{max}}{}^{-(M-3)}\bigl(1-h_{M}(0.45)\bigr)-h_{M}(\gamma_{\mathrm{max}})S_{M-2}(\gamma_{\mathrm{max}})}{S_{M-2}(\gamma_{\mathrm{max}})+\gamma_{\mathrm{max}}+\frac{1}{0.45}\gamma_{\mathrm{max}}{}^{-(M-3)}}.$
Taking the above two displays together, we reach
$u_{1}+\phi_{M}(u_{M-1})$ $>(1+\gamma_{\mathrm{min}})\left\{\frac{\frac{1}{0.45}\gamma_{\mathrm{max}}{}^{-(M-3)}\bigl(1-h_{M}(0.45)\bigr)-h_{M}(\gamma_{\mathrm{max}})S_{M-2}(\gamma_{\mathrm{max}})}{S_{M-2}(\gamma_{\mathrm{max}})+\gamma_{\mathrm{max}}+\frac{1}{0.45}\gamma_{\mathrm{max}}{}^{-(M-3)}}\right\}$
$\quad+(h_{M}(\frac{1}{2})-\frac{1}{2}h_{M}^{\prime}(\frac{1}{2}))\gamma_{\mathrm{min}}+h_{M}(\gamma_{\mathrm{min}}).$
The RHS exceeds 1 and one has $u_{1}+\phi_{M}(u_{M-1})>1$, which contradicts with $u_{M}=u_{1}+\phi_{M}(u_{M-1})=1$; see Figure 6. As a result, we must have $\gamma_{M-1}\geq 0.5$.
By the second item of Proposition 3, we know that $\gamma_{i}$ is decreasing and hence $\gamma_{i}\geq\gamma_{M-1}\geq 0.5$ for $i\leq M-2$. Since $\bm{u}$ lies in the interior of $\mathcal{U}_{M}$, we have $u_{i-1}<u_{i}$, and this leads to
$u_{i-1}\leq u_{i}\cdot(\gamma_{i+1}+\frac{1}{2}),\quad\forall 2\leq i\leq M-2.$
In other words, we have established the desired inequalities except for the last one. The remaining of the proof is devoted to show $u_{M-2}<u_{M-1}(\gamma_{M}+\frac{1}{2}).$
Step 2: Establishing $u_{M-2}<u_{M-1}(\gamma_{M}+\frac{1}{2})$.
We split the proof into two cases: (1) $M\leq 6$, and $M\geq 7$.
A key lemma is the following.
Lemma 17.
Assume $b<\gamma_{M}<1/2$ for some $b\in(\gamma_{\mathrm{min}},1/2)$. If
$\gamma_{M-1}<\frac{(1+h_{M}^{\prime}(b))(1+\gamma_{M}(\frac{1}{2}-b))}{\frac{1}{2}-b},$ (41)
then $u_{M-2}<u_{M-1}(\gamma_{M}+\frac{1}{2}).$
When $M\leq 6$, setting $b=0.45$, we have
$\frac{(1+h_{M}^{\prime}(b))(1+\gamma_{M}(\frac{1}{2}-b))}{\frac{1}{2}-b}>1.$
Since $\gamma_{M-1}<1$, the condition (41) holds. We can then apply Lemma 17 to reach the conclusion that $u_{M-2}<u_{M-1}(\gamma_{M}+\frac{1}{2})$.
In the case when $M\geq 7$, we further consider two subcases: (1) $d\geq 2$, and (2) $d=1$.
When $d\geq 2$, we prove that $\gamma_{M}\geq 1/2$. To see this, suppose that $\gamma_{M}<1/2$, we apply Lemma 15 with $c=1/2$ to obtain $u_{1}+\phi_{M}(u_{M-1})>1$, which is a contradiction with $u_{M}=u_{1}+\phi_{M}(u_{M-1})=1$. Hence we have $\gamma_{M}\geq 1/2$, that further implies $u_{M-2}<u_{M-1}(\gamma_{M}+\frac{1}{2})$.
For $d=1$, there are again two cases. When $M\geq 10$, we can again prove $\gamma_{M}\geq 1/2$ by Lemma 15. When $M=7,8,9$, Lemma 15 gives $\gamma_{M}\geq 0.47,0.48,0.49,$ respectively. Plugging in the corresponding lower bound value of $\gamma_{M}$ as $b$ to Lemma 17, one can verify that the precondition holds and therefore $u_{M-2}<u_{M-1}(\gamma_{M}+\frac{1}{2}).$
Figure 6: Lower bound of $u_{M}$ vs. batch budget $M$ when $\gamma_{M-1}<0.5$.
B.3.3 Remaining proofs
Proof of Lemma 15.
Recall that $\gamma_{M}=\operatorname*{arg\,min}_{\gamma\in\mathcal{S}}\gamma u_{M-1}+h_{M}(\gamma)$. Since $\gamma_{M}<c\leq 1/2$, one has either $u_{M-1}+h_{M}^{\prime}(\gamma_{M})=0$ or $\gamma_{M}=\gamma_{\mathrm{min}}$. In both cases, convexity implies $u_{M-1}\geq-h_{M}^{\prime}(\gamma_{M})$. In addition, we know that the function $h_{M}^{\prime}(\cdot)$ is strictly increasing and $\gamma_{M}<c$, we have $u_{M-1}\geq-h_{M}^{\prime}(\gamma_{M})>-h_{M}^{\prime}(c)$. Consequently,
$u_{1}+\min_{\gamma\in\mathcal{S}}\gamma u_{M-1}+h_{M}(\gamma)$ $>u_{1}+\min_{\gamma\in\mathcal{S}}\gamma(-h_{M}^{\prime}(c))+h_{M}(\gamma)$
$=u_{1}-ch_{M}^{\prime}(c)+h_{M}(c)$
$\geq\frac{c^{-1}\gamma_{\mathrm{max}}{}^{-(M-3)}\bigl(1-h_{M}(c)\bigr)-h_{M}(\gamma_{\mathrm{max}})S_{M-2}(\gamma_{\mathrm{max}})}{S_{M-2}(\gamma_{\mathrm{max}})+\gamma_{\mathrm{max}}+c^{-1}\gamma_{\mathrm{max}}{}^{-(M-3)},}-ch_{M}^{\prime}(c)+h_{M}(c),$
where the second step is due to $\operatorname*{arg\,min}_{\gamma\in\mathcal{S}}-\gamma h_{M}^{\prime}(c)+h_{M}(c)=c$, and the last step applies Lemma 11. ∎
Proof of Lemma 16.
Recall that $\gamma_{M-1}=\operatorname*{arg\,min}_{\gamma\in\mathcal{S}}\gamma u_{M-2}+h_{M}(\gamma)$. Since by assumption $\gamma_{M-1}<1/2$, one has either $u_{M-2}+h_{M}^{\prime}(\gamma_{M-1})=0$ or $\gamma_{M}=\gamma_{\mathrm{min}}$. In both cases, convexity implies $u_{M-2}\geq-h_{M}^{\prime}(\gamma_{M-1}).$ Since the function $h_{M}^{\prime}(\cdot)$ is strictly increasing and $\gamma_{M-1}<1/2$, we have $u_{M-1}\geq-h_{M}^{\prime}(\gamma_{M-1})>-h_{M}^{\prime}(1/2)$. Consequently,
$u_{1}+\phi_{M}(u_{M-1})$ $=u_{1}+\phi_{M}(u_{1}+\phi_{M}(u_{M-2}))$
$\geq u_{1}+\phi_{M}(u_{1}+\min_{\gamma\in\mathcal{S}}-h_{M}^{\prime}(1/2)\gamma+h_{M}(\gamma))$
$\overset{}{=}u_{1}+\phi_{M}(u_{1}-\frac{1}{2}h_{M}^{\prime}(\frac{1}{2})+h_{M}(\frac{1}{2}))$
$=u_{1}+\min_{\gamma\in\mathcal{S}}(u_{1}-\frac{1}{2}h_{M}^{\prime}(\frac{1}{2})+h_{M}(\frac{1}{2}))\gamma+h_{M}(\gamma),$
where the penultimate step is due to $\operatorname*{arg\,min}_{\gamma\in\mathcal{S}}-\gamma h_{M}^{\prime}(1/2)+h_{M}(1/2)=1/2$. By Lemma 11 with $c=0.45$, one has $u_{1}-\frac{1}{2}h_{M}^{\prime}(\frac{1}{2})+h_{M}(\frac{1}{2})>-h_{M}^{\prime}(\gamma_{\mathrm{min}})$. Hence, the minimizer of the above function is attained at $\gamma=\gamma_{\mathrm{min}}$. We reach
$u_{1}+\phi_{M}(u_{M-1})$ $\geq u_{1}+\min_{\gamma\in\mathcal{S}}(u_{1}-\frac{1}{2}h_{M}^{\prime}(\frac{1}{2})+h_{M}(\frac{1}{2}))\gamma+h_{M}(\gamma)$
$\overset{}{=}u_{1}+(u_{1}-\frac{1}{2}h_{M}^{\prime}(\frac{1}{2})+h_{M}(\frac{1}{2}))\gamma_{\mathrm{min}}+h_{M}(\gamma_{\mathrm{min}})$
$=(1+\gamma_{\mathrm{min}})u_{1}+(-\frac{1}{2}h_{M}^{\prime}(\frac{1}{2})+h_{M}(\frac{1}{2}))\gamma_{\mathrm{min}}+h_{M}(\gamma_{\mathrm{min}}).$
This completes the proof. ∎
Proof of Lemma 17.
Denote by $\Delta_{i}=u_{i+1}-u_{i}$. It suffices to show $u_{M-1}(1/2-\gamma_{M})<\Delta_{M-2}$. Since
$u_{M-1}(1/2-\gamma_{M})$ $=(1-\Delta_{M-1})(1/2-\gamma_{M})\leq(1-\gamma_{M}\Delta_{M-2})(1/2-\gamma_{M}),$
where the inequality is due to Lemma 14 and the assumption that $\gamma_{M}<1/2$. It boils down to establishing $(1-\gamma_{M}\Delta_{M-2})(1/2-\gamma_{M})<\Delta_{M-2}$, which is equivalent to showing that
$\Delta_{M-2}>\frac{\frac{1}{2}-\gamma_{M}}{1+\gamma_{M}(\frac{1}{2}-\gamma_{M})}.$
Note that
$\Delta_{M-2}=\frac{\Delta_{M-2}}{\Delta_{M-1}}\cdot\Delta_{M-1}\geq\frac{1}{\gamma_{M-1}}\cdot\Delta_{M-1}=\frac{1}{\gamma_{M-1}}\cdot(1-u_{M-1}),$
where the first inequality again uses Lemma 14. Further note that since $\gamma_{M}\in(\gamma_{\mathrm{min}},1/2)$ minimizes $u_{M-1}g+h_{M}(g)$, we obtain $u_{M-1}=-h_{M}^{\prime}(\gamma_{M})$. This allows us to lower bound $\Delta_{M-2}$ as
$\Delta_{M-2}\geq\frac{1}{\gamma_{M-1}}\cdot(1+h_{M}^{\prime}(\gamma_{M})).$
In all, it suffices to show that
$\gamma_{M-1}<\frac{(1+h_{M}^{\prime}(b))(1+\gamma_{M}(\frac{1}{2}-b))}{\frac{1}{2}-b}$ (42)
Under the assumption (41), we have
$\gamma_{M-1}<\frac{(1+h_{M}^{\prime}(b))(1+\gamma_{M}(\frac{1}{2}-b))}{\frac{1}{2}-b}.$
Note that the RHS as a function of $b$ is increasing when $b\in(\gamma_{\mathrm{min}},1/2)$, we then have
$\gamma_{M-1}<\frac{(1+h_{M}^{\prime}(b))(1+\gamma_{M}(\frac{1}{2}-b))}{\frac{1}{2}-b}<\frac{(1+h_{M}^{\prime}(\gamma_{M}))(1+\gamma_{M}(\frac{1}{2}-\gamma_{M}))}{\frac{1}{2}-\gamma_{M}}.$
This is equivalent to our final goal (42). Hence the proof is completed. ∎
Appendix C Remaining proofs for the lower bound
C.1 Remaining proofs of Section 5.2
Proof of Proposition 6.
It is straightforward to check $f^{(1)}_{S,\sigma,i}$ satisfies the smoothness condition.
We now verify the margin condition. If $\delta<\delta_{i}$, by design we have
$P_{X}\left(0<\left|f^{(1)}_{S,\sigma,i}(X)-\frac{1}{2}\right|\leq\delta\right)=0.$
Otherwise, choose $\ell\leq i$ s.t. $\delta_{\ell}\leq\delta<\delta_{\ell-1}$. Then $\{0<|f^{(1)}-\tfrac{1}{2}|\leq\delta\}\subset\bigcup_{m\geq\ell}\bigcup_{j\in S_{m}}B_{m,j}$. As a result, we have
$\mathbb{P}(0<|f^{(1)}-\tfrac{1}{2}|\leq\delta)\ \leq\sum_{m\geq\ell}^{i}\sum_{j\in S_{m}}\mathbb{P}(B_{m,j})=\sum_{m\geq\ell}^{i}s_{m}(Mz_{m})^{-d},$
where the last relation arises from the covariate distribution (16).
Recall that $s_{m}=M^{-1}(Mz_{m})^{d-\alpha_{m}\beta}$. We further have
$\mathbb{P}(0<|f^{(1)}-\tfrac{1}{2}|\leq\delta)\ \leq\sum_{m\geq\ell}^{i}s_{m}(Mz_{m})^{-d}=\sum_{m\geq\ell}^{i}M^{-1}(Mz_{m})^{-\alpha_{m}\beta}.$
Since $(Mz_{m})^{-\alpha_{m}\beta}\leq(Mz_{l})^{-\alpha_{m}\beta}$ for $m\geq l$ and $\alpha_{m}\geq\alpha_{i}$ for for $m\leq i$, we arrive at
$\mathbb{P}(0<|f^{(1)}-\tfrac{1}{2}|\leq\delta)\ \leq(Mz_{l})^{-\alpha_{i}\beta}\leq(\frac{\delta}{D_{\phi}})^{\alpha_{i}}.$
We therefore established the smoothness and the margin condition. ∎
C.2 Remaining proofs of Section 5.3
Proof of Lemma 2.
It suffices to bound their KL-divergence. We can compute
$\mathrm{KL}(\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{n},\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=1}^{n})$ $\overset{\mathrm{(k)}}{\leq}8\mathbb{E}_{\Gamma,\pi;\sigma_{i,j}=-1}[\sum_{t=1}^{n}(f_{\sigma_{i,j}=-1}(X_{t})-f_{\sigma_{i,j}=1}(X_{t}))^{2}\mathbf{1}\{\pi_{t}(X_{t})=1\}]$
$\overset{\mathrm{(ii)}}{\leq}32D_{\phi}^{2}(Mz_{i})^{-2\beta}\mathbb{E}_{\Gamma,\pi;\sigma_{i,j}=-1}[\sum_{t=1}^{n}\mathbf{1}\{\pi_{t}(X_{t})=1,X_{t}\in B_{i,j}\}]$
$\overset{\mathrm{(iii)}}{=}32D_{\phi}^{2}(Mz_{i})^{-(2\beta+d)}\sum_{t=1}^{n}\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{t}(\pi_{t}(X_{t})=1\mid X_{t}\in B_{i,j})$
$\overset{\mathrm{(iv)}}{\leq}32D_{\phi}^{2}(Mz_{i})^{-(2\beta+d)}n\leq 2n(Mz_{i})^{-(2\beta+d)}.$
Here, step (k) uses the standard decomposition of KL divergence and Bernoulli reward structure; step (ii) is due to the definition of $f_{\omega}$; step (iii) uses $\mathbb{P}(X_{t}\in B_{i,j})=1/(Mz_{i})^{d}$, and step (iv) arises from $\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{t}(\pi_{t}(X_{t})=1\mid X_{t}\in B_{i,j})\leq 1$ for any $1\leq t\leq n$.
By Pinsker’s inequality,
$\|\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{n}-\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=1}^{n}\|_{\mathrm{TV}}\leq\sqrt{\frac{1}{2}\mathrm{KL}(\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=-1}^{n},\mathbb{P}_{\Gamma,\pi;\sigma_{i,j}=1}^{n})}\leq\sqrt{n(Mz_{i})^{-(2\beta+d)}}.$
This finishes the proof. ∎
C.3 Remaining proofs of Section 5.4
Proof of Lemma 6.
Throughout the proof, we drop the subscript on $i$. Recall
$r(\sigma)\coloneqq\frac{(\tfrac{1}{2}+\sigma\delta)^{R}(\tfrac{1}{2}-\sigma\delta)^{N-R}}{(\tfrac{1}{2})^{N}}=(1+2\sigma\delta)^{R}(1-2\sigma\delta)^{N-R},\quad\sigma\in\{+1,-1\},$
and
$m\coloneqq\tfrac{1}{2}\big(r(+1)+r(-1)\big).$
We want to compute $\mathbb{E}_{0}[m\mid N]$ and $\mathbb{E}_{0}[m^{2}\mid N]$. To start with,
$\mathbb{E}_{0}[r(\sigma)\mid N]$ $=\sum_{x=0}^{N}\binom{N}{x}2^{-N}\,(1+2\sigma\delta)^{x}(1-2\sigma\delta)^{N-x}$
$=2^{-N}\sum_{x=0}^{N}\binom{N}{x}\big(1+2\sigma\delta\big)^{x}\big(1-2\sigma\delta\big)^{N-x}$
$=2^{-N}\,\big[(1+2\sigma\delta)+(1-2\sigma\delta)\big]^{N}\quad\text{(binomial theorem)}$
$=2^{-N}\,(2)^{N}=1.$
Therefore,
$\mathbb{E}_{0}[m\mid N]=\tfrac{1}{2}\big(\mathbb{E}_{0}[r(+1)\mid N]+\mathbb{E}_{0}[r(-1)\mid N]\big)=1.$
Next, we deal with the second moment. Expanding $m^{2}$,
$m^{2}=\tfrac{1}{4}\Big(r(+1)^{2}+2\,r(+1)r(-1)+r(-1)^{2}\Big).$
We will evaluate the three expectations separately.
$\mathbb{E}_{0}[r(+1)^{2}\mid N]$ $=\sum_{x=0}^{N}\binom{N}{x}2^{-N}(1+2\delta)^{2x}(1-2\delta)^{2(N-x)}$
$=2^{-N}\big[(1+2\delta)^{2}+(1-2\delta)^{2}\big]^{N}$
$=2^{-N}\big(2+8\delta^{2}\big)^{N}=(1+4\delta^{2})^{N}.$
By symmetry, $\mathbb{E}_{0}[r(-1)^{2}\mid N]=(1+4\delta^{2})^{N}$. For the cross-product,
$r(+1)r(-1)$ $=\big[(1+2\delta)^{R}(1-2\delta)^{N-R}\big]\big[(1-2\delta)^{R}(1+2\delta)^{N-R}\big]$
$=\big((1+2\delta)(1-2\delta)\big)^{R}\big((1-2\delta)(1+2\delta)\big)^{N-R}$
$=(1-4\delta^{2})^{R}(1-4\delta^{2})^{N-R}=(1-4\delta^{2})^{N},$
which is constant in $R$, hence
$\mathbb{E}_{0}[r(+1)r(-1)\mid N]=(1-4\delta^{2})^{N}.$
Putting things together,
$\mathbb{E}_{0}[m^{2}\mid N]$ $=\tfrac{1}{4}\Big(\mathbb{E}_{0}[r(+1)^{2}\mid N]+2\,\mathbb{E}_{0}[r(+1)r(-1)\mid N]+\mathbb{E}_{0}[r(-1)^{2}\mid N]\Big)$
$=\tfrac{1}{4}\Big((1+4\delta^{2})^{N}+2(1-4\delta^{2})^{N}+(1+4\delta^{2})^{N}\Big)$
$=\tfrac{1}{2}\Big((1+4\delta^{2})^{N}+(1-4\delta^{2})^{N}\Big).$
This finishes the proof. ∎
Proof of Lemma 7.
By the law of total expectation,
$\mathbb{E}[t^{L}]$ $=\mathbb{E}[\mathbb{E}[t^{L}\mid S]]$
$=\mathbb{E}[\mathbb{E}[t^{\sum_{i\in S}\mathbf{1}\{i\in S^{\prime}\}}\mid S]]$
$\leq\mathbb{E}[\mathbb{E}[t^{\mathrm{Binomial}(s,\frac{s}{z})}\mid S]],$
where the last step applies Lemma 1.1 in Bardenet and Maillard (2015). Using the PGF of binomial distribution, we obtain
$\mathbb{E}[t^{L}]$ $\leq\mathbb{E}[\mathbb{E}[t^{\mathrm{Binomial}(s,\frac{s}{z})}\mid S]]$
$=\left(1+\frac{s}{z}(t-1)\right)^{s}\leq\exp\left(\frac{s^{2}}{z}(t-1)\right).$
This completes the proof. ∎
Appendix D Proof of the upper bound
This section is devoted to establishing Lemma 1, whose proof follows the framework developed in Perchet and Rigollet (2013); Jiang and Ma (2025).
To start with, recall $\mathcal{T}$ is a tree of depth $M$, whose root (depth 0) represents the whole covariate space $\mathcal{X}$. The tree is recursively defined as the following: for any $i\geq 1$, each node at depth i1i-1 is split into $g_{i-1}^{d}$ children. Consequently, a node at depth $i$ has width $w_{i}=g_{i-1}^{-1}\cdot w_{i-1}=(\prod_{l=0}^{i-1}g_{l})^{-1}$. For any bin $C\in\mathcal{T}$, denote its parent by $\mathsf{p}(C)=\{C^{\prime}\in\mathcal{T}:C\in\mathsf{child}(C^{\prime})\}$. Define $\mathsf{p}^{1}(C)=\mathsf{p}(C)$ and $\mathsf{p}^{k}(C)=\mathsf{p}(\mathsf{p}^{k-1}(C))$ for $k\geq 2$. Let $\mathcal{P}(C)=\{C^{\prime}\in\mathcal{T}:C^{\prime}=\mathsf{p}^{k}(C)\textrm{ for some }k\geq 1\}$ be the set of ancestors of the bin $C$. Denote by $\mathcal{L}_{0}=\{\mathcal{X}\}$, and let $\mathcal{L}_{t}$ be the set of active bins at time $t$. It is easy to see $\mathcal{L}_{t}=\mathcal{B}_{1}$ for $1\leq t\leq t_{1}$, where $\mathcal{B}_{1}$ are all the bins in the first layer.
D.1 Introducing the good events
Fix a batch $i\geq 1$, for any $C\in\mathcal{L}_{t_{i-1}+1}$, define
$m_{C,i}\coloneqq\sum_{t=t_{i-1}+1}^{t_{i}}\mathbf{1}\{X_{t}\in C\},$
which is the number of times the covariates land into bin $C$ during batch $i$. The expectation of $m_{C,i}$ is equal to
$m_{C,i}^{\star}=\mathbb{E}[m_{C,i}]=(t_{i}-t_{i-1})P_{X}(X\in C).$
Since $\bm{u}$ lies in the interior of $\mathcal{U}_{M}$ and the split factors satisfy equation (11), we have $|C|=w_{i}=(\prod_{l=0}^{i-1}g_{l})^{-1}\geq T^{-1/(2\beta+d)}.$ The lemma below says $m_{C,i}$ stays closely to its expectation $m_{C,i}^{\star}$ for all $C\in\mathcal{T}$.
Lemma 18.
Assume that $M\leq D_{1}\log(T)$ for some constant $D_{1}>0$. With probability at least 11/T1-1/T, for all $1\leq i\leq M$ and $C\in\mathcal{L}_{t_{i-1}+1}$, one has
$\frac{1}{2}m_{C,i}^{\star}\leq m_{C,i}\leq\frac{3}{2}m_{C,i}^{\star}.$
Proof of Lemma 18
Fix the batch index $i$, and a node $C$ in layer-$i$ of the tree $\mathcal{T}$. If $P_{X}(C)=0$, then $m_{C,i}=m_{C,i}^{\star}=0$ almost surely. For the remaining part of the proof, we assume $P_{X}(C)>0$. By relation (12), we have
$m_{C,i}^{\star}$ $=(t_{i}-t_{i-1})P_{X}(X\in C)$
$\asymp|C|^{-(2\beta+d)}\log(T|C|^{d})P_{X}(X\in C)$
$\overset{\mathrm{(i)}}{\apprge}|C|^{-2\beta}\overset{\mathrm{}}{\geq}g_{0}^{2\beta},$
where step (i) uses Assumption 1. Since $g_{0}=\lfloor T^{\frac{1}{2\beta+d}\cdot u_{1}}\rfloor$ and $u_{1}>0$, we reach $m_{C,i}^{\star}\geq\frac{3}{4}\log(2T^{2})$ for all $i$ and $C$, as long as $T$ is sufficiently large. This allows us to invoke Chernoff’s bound to obtain that with probability at most $1/T^{2}$,
$\left|\sum\nolimits_{t=t_{i-1}+1}^{t_{i}}\mathbf{1}\{X_{t}\in C\}-m_{C,i}^{\star}\right|\geq\sqrt{3\log(2T^{2})m_{C,i}^{\star}}.$
Denote $E^{c}=\{\exists 1\leq i\leq M,C\in\mathcal{L}_{t_{i-1}+1}\text{ such that }\mid\sum_{t=t_{i-1}+1}^{t_{i}}\mathbf{1}\{X_{t}\in C\}-m_{C,i}^{\star}\mid\geq\sqrt{3\log(2T^{2})m_{C,i}^{\star}}\}$. Applying union bound to reach
$\mathbb{P}(E^{c})$ $\leq\sum_{C\in\mathcal{T}}\frac{1}{T^{2}}\overset{\mathrm{(ii)}}{\leq}\frac{1}{T^{2}}\left(\sum_{i=1}^{M}(\prod_{l=0}^{i-1}g_{l})^{d}\right)\overset{\mathrm{(iii)}}{\leq}\frac{1}{T^{2}}\cdot M\cdot(\prod_{l=0}^{M-1}g_{l})^{d},$
where step (ii) sums over all possible nodes of $\mathcal{T}$ across batches, and step (iii) is due to $(\prod_{l=0}^{i-1}g_{l})^{d}\leq(\prod_{l=0}^{M-1}g_{l})^{d}$ for any $1\leq i\leq M$. Since $g_{M-1}=1$, we further obtain
$\mathbb{P}(E^{c})$ $\leq\frac{1}{T^{2}}\cdot M\cdot(\prod_{l=0}^{M-2}g_{l})^{d}\overset{\mathrm{(iv)}}{\leq}\frac{1}{T^{2}}\cdot M\cdot t_{M-1}^{\frac{d}{2\beta+d}}\overset{\mathrm{}}{\leq}D_{1}\frac{1}{T^{2}}\cdot T^{\frac{d}{2\beta+d}}\leq\frac{1}{T},$
where step (iv) invokes relation (12). This completes the proof.
Denote the above event by $E$. By assumption $M\leq D_{1}\log(T)$, we use Lemma 18 to reach
$\mathbb{E}[R_{T}(\hat{\Gamma}_{\bm{u}},\hat{\pi}_{\bm{u}})\mathbf{1}(E^{c})]\leq T\mathbb{P}(E^{c})=1,$
which means the regret incurred when $E$ does not happen is negligible. For the remaining proof, the task becomes controlling $\mathbb{E}[R_{T}(\hat{\pi})\mathbf{1}(E)]$.
Next, we turn to the arm elimination part. For each bin $C\in\mathcal{L}_{t_{i}}$, denote by $\mathcal{I}_{C}^{\prime}$ the set of remaining arms at the end of batch $i$, i.e., after Algorithm 2 is invoked. Define
$\underline{\mathcal{I}}_{C}=\left\{k\in\{1,-1\}:\sup_{x\in C}f^{\star}(x)-f^{(k)}(x)\leq c_{0}|C|^{\beta}\right\},$
$\bar{\mathcal{I}}_{C}$ $=\left\{k\in\{1,-1\}:\sup_{x\in C}f^{\star}(x)-f^{(k)}(x)\leq c_{1}|C|^{\beta}\right\},$
where $c_{0}=2Ld^{\beta/2}+1$ and $c_{1}=8c_{0}$. By definition,
$\underline{\mathcal{I}}_{C}\subseteq\bar{\mathcal{I}}_{C}.$
Define the event $\mathcal{A}_{C}=\{\underline{\mathcal{I}}_{C}\subseteq\mathcal{I}_{C}^{\prime}\subseteq\bar{\mathcal{I}}_{C}\}$. Besides, define $\mathcal{G}_{C}=\cap_{C^{\prime}\in\mathcal{P}(C)}\mathcal{A}_{C^{\prime}}$. For $i\geq 1$, recall that $\mathcal{B}_{i}$ is the collection of bins $C$ with $|C|=(\prod_{l=0}^{i-1}g_{l})^{-1}=w_{i}$. The following lemma adapted from Jiang and Ma (2025) shows successive elimination succeeds with high probability.
Lemma 19.
For any $1\leq i\leq M-1$ and $C\in\mathcal{B}_{i}$ such that $P_{X}(C)>0$, we have
$\mathbb{P}(E\cap\mathcal{G}_{C}\cap\mathcal{A}_{C}^{c})\leq\frac{4m_{C,i}^{\star}}{T|C|^{d}}.$
D.2 Regret decomposition
For any bin $C\in\mathcal{T}$, we consider the following two sources of regret incurred on it. First, define
$r_{T}^{\textrm{live}}(C)\coloneqq\sum_{t=1}^{T}\left(f^{\star}(X_{t})-f_{\pi_{t}(X_{t})}(X_{t})\right)\mathbf{1}(X_{t}\in C)\mathbf{1}(C\in\mathcal{L}_{t}).$
Besides, denote by $\mathcal{J}_{t}\coloneqq\cup_{s\leq t}\mathcal{L}_{s}$ the set of bins that have been live up to time $t$. Define
$r_{T}^{\textrm{born}}(C)\coloneqq\sum_{t=1}^{T}\left(f^{\star}(X_{t})-f_{\pi_{t}(X_{t})}(X_{t})\right)\mathbf{1}(X_{t}\in C)\mathbf{1}(C\in\mathcal{J}_{t}).$
Due to the structure of the tree $\mathcal{T}$, we have
$r_{T}^{\textrm{born}}(C)$ $=r_{T}^{\textrm{live}}(C)+\sum_{C^{\prime}\in\mathsf{child}(C)}r_{T}^{\textrm{born}}(C^{\prime})$
$=r_{T}^{\textrm{born}}(C)\mathbf{1}(\mathcal{A}_{C}^{c})+r_{T}^{\textrm{live}}(C)\mathbf{1}(\mathcal{A}_{C})+\sum_{C^{\prime}\in\mathsf{child}(C)}r_{T}^{\textrm{born}}(C^{\prime})\mathbf{1}(\mathcal{A}_{C}).$
The following regret decomposition is an immediate consequence of iteratively applying the relation above to each level of the tree.
$R_{T}(\hat{\Gamma}_{\bm{u}},\hat{\pi}_{\bm{u}})$ $=\mathbb{E}[r_{T}^{\textrm{born}}(\mathcal{X})]$
$=\sum_{C^{\prime}\in\mathsf{child}(\mathcal{X})}\mathbb{E}[r_{T}^{\textrm{born}}(C^{\prime})]$
$=\sum_{1\leq i\leq M-1}\left(\underbrace{\sum_{C\in\mathcal{B}_{i}}\mathbb{E}[r_{T}^{\textrm{born}}(C)\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C}^{c})]}_{\eqqcolon U_{i}}+\underbrace{\sum_{C\in\mathcal{B}_{i}}\mathbb{E}[r_{T}^{\textrm{live}}(C)\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C})]}_{\eqqcolon V_{i}}\right)$
$\quad+\sum_{C\in\mathcal{B}_{M}}\mathbb{E}[r_{T}^{\textrm{live}}(C)\mathbf{1}(\mathcal{G}_{C})],$
where the second step is due to $r_{T}^{\textrm{live}}(\mathcal{X})=0$ (note $\mathcal{X}\notin\mathcal{L}_{t}$ for any $1\leq t\leq T$). Now that we have a decomposition of the regret, the task becomes bounding $V_{i},U_{i}$ and the regret of the last batch separately.
We first consider the case of $\alpha\leq d/\beta$.
Upper bounding term $V_{i}$.
Fix some $1\leq i\leq M-1$, and some bin $C\in\mathcal{B}_{i}$. The event $\mathcal{G}_{C}$ implies $\mathcal{I}_{\mathsf{p}(C)}^{{}^{\prime}}\subseteq\bar{\mathcal{I}}_{\mathsf{p}(C)}$. Namely, for any $k\in\mathcal{I}_{\mathsf{p}(C)}^{{}^{\prime}}$,
$\sup_{x\in\mathsf{p}(C)}f^{\star}(x)-f^{(k)}(x)\leq c_{1}|\mathsf{p}(C)|^{\beta}.$
Consequently, for any $x\in C$, and $k\in\mathcal{I}_{\mathsf{p}(C)}^{{}^{\prime}}$,
$\left(f^{\star}(x)-f^{(k)}(x)\right)\bm{1}\{\mathcal{G}_{C}\}\leq c_{1}|\mathsf{p}(C)|^{\beta}\mathbf{1}(0<\left|f_{1}(x)-f^{(-1)}(x)\right|\leq c_{1}|\mathsf{p}(C)|^{\beta}).$ (43)
This leads to
$\mathbb{E}[r_{T}^{\textrm{live}}(C)\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C})]$
$\quad\overset{\mathrm{}}{\leq}\mathbb{E}\left[\sum_{t=1}^{T}c_{1}|\mathsf{p}(C)|^{\beta}\mathbf{1}(0<\left|f_{1}(X_{t})-f^{(-1)}(X_{t})\right|\leq c_{1}|\mathsf{p}(C)|^{\beta})\mathbf{1}(X_{t}\in C,C\in\mathcal{L}_{t})\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C})\right]$
$\quad\overset{\mathrm{(i)}}{\leq}c_{1}|\mathsf{p}(C)|^{\beta}\mathbb{E}\left[\sum_{t=t_{i-1}+1}^{t_{i}}\mathbf{1}(0<\left|f_{1}(X_{t})-f^{(-1)}(X_{t})\right|\leq c_{1}|\mathsf{p}(C)|^{\beta},X_{t}\in C)\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C})\right]$
$\quad\overset{\mathrm{(ii)}}{\leq}c_{1}|\mathsf{p}(C)|^{\beta}\sum_{t=t_{i-1}+1}^{t_{i}}\mathbb{P}(0<\left|f_{1}(X_{t})-f^{(-1)}(X_{t})\right|\leq c_{1}|\mathsf{p}(C)|^{\beta},X_{t}\in C)$
$\quad=c_{1}|\mathsf{p}(C)|^{\beta}(t_{i}-t_{i-1})\mathbb{P}(0<\left|f_{1}(X)-f^{(-1)}(X)\right|\leq c_{1}|\mathsf{p}(C)|^{\beta},X\in C).$
Here, step (i) can be deduced from considering the cases of whether $C$ is split or not; step (ii) is because $\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C})\leq 1$.
Summing over all bins in $\mathcal{B}_{i}$, we reach
$\sum_{C\in\mathcal{B}_{i}}\mathbb{E}[r_{T}^{\textrm{live}}(C)\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C})]$ $\leq\sum_{C\in\mathcal{B}_{i}}c_{1}w_{i-1}^{\beta}(t_{i}-t_{i-1})\mathbb{P}(0<\left|f_{1}(X)-f^{(-1)}(X)\right|\leq c_{1}|\mathsf{p}(C)|^{\beta},X\in C)$
$=c_{1}w_{i-1}^{\beta}(t_{i}-t_{i-1})\sum_{C\in\mathcal{B}_{i}}\mathbb{P}(0<\left|f_{1}(X)-f^{(-1)}(X)\right|\leq c_{1}w_{i-1}^{\beta},X\in C)$
$=c_{1}w_{i-1}^{\beta}(t_{i}-t_{i-1})\mathbb{P}(0<\left|f_{1}(X)-f^{(-1)}(X)\right|\leq c_{1}w_{i-1}^{\beta}),$ (44)
where the penultimate step is due to $|\mathsf{p}(C)|=w_{i-1}$. We can apply the margin condition to obtain
$V_{i}=\sum_{C\in\mathcal{B}_{i}}\mathbb{E}[r_{T}^{\textrm{live}}(C)\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C})]$ $\leq(t_{i}-t_{i-1})\cdot[c_{1}w_{i-1}^{\beta}]^{1+\alpha}\cdot D_{0}.$
Upper bounding term $U_{i}$.
Fix some $1\leq i\leq M-1$, and some bin $C\in\mathcal{B}_{i}$ such that $P_{X}(C)>0$. By relation (43),
$\mathbb{E}$ $[r_{T}^{\textrm{born}}(C)\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C}^{c})]$
$\leq\mathbb{E}\left[\sum_{t=1}^{T}c_{1}|\mathsf{p}(C)|^{\beta}\mathbf{1}(0<\left|f_{1}(X_{t})-f^{(-1)}(X_{t})\right|\leq c_{1}|\mathsf{p}(C)|^{\beta})\mathbf{1}(X_{t}\in C,C\in\mathcal{J}_{t})\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C}^{c})\right]$
$\leq c_{1}|\mathsf{p}(C)|^{\beta}T\mathbb{P}(0<\left|f_{1}(X)-f^{(-1)}(X)\right|\leq c_{1}|\mathsf{p}(C)|^{\beta},X\in C)\mathbb{P}(\mathcal{G}_{C}\cap\mathcal{A}_{C}^{c})$
$\leq c_{1}|\mathsf{p}(C)|^{\beta}T\mathbb{P}(0<\left|f_{1}(X)-f^{(-1)}(X)\right|\leq c_{1}|\mathsf{p}(C)|^{\beta},X\in C)\frac{4m_{C,i}^{\star}}{T|C|^{d}}$
$\leq 4\bar{c}c_{1}w_{i-1}^{\beta}\mathbb{P}(0<\left|f_{1}(X)-f^{(-1)}(X)\right|\leq c_{1}w_{i-1}^{\beta},X\in C)(t_{i}-t_{i-1}),$
where the penultimate step uses Lemma 19, and the last step is due to Assumption 1. Consequently, we reach
$U_{i}$ $=\sum_{C\in\mathcal{B}_{i}}\mathbb{E}[r_{T}^{\textrm{born}}(C)\mathbf{1}(\mathcal{G}_{C}\cap\mathcal{A}_{C}^{c})]$
$\leq 4\bar{c}c_{1}w_{i-1}^{\beta}(t_{i}-t_{i-1})\sum_{C\in\mathcal{B}_{i}}\mathbb{P}(0<\left|f_{1}(X)-f^{(-1)}(X)\right|\leq c_{1}w_{i-1}^{\beta},X\in C)$
$=4\bar{c}c_{1}w_{i-1}^{\beta}(t_{i}-t_{i-1})\mathbb{P}(0<\left|f_{1}(X)-f^{(-1)}(X)\right|\leq c_{1}w_{i-1}^{\beta}),$ (45)
By the margin condition, we get
$U_{i}\leq 4D_{0}\bar{c}(t_{i}-t_{i-1})[c_{1}w_{i-1}^{\beta}]^{1+\alpha}.$
Regret of last Batch.
For $C\in\mathcal{B}_{M}$, similarly we have
$\mathbb{E}[r_{T}^{\textrm{live}}(C)\mathbf{1}(\mathcal{G}_{C})]$ $\leq c_{1}|\mathsf{p}(C)|^{\beta}(T-t_{M-1})\mathbb{P}(0<\left|f_{1}(X)-f^{(-1)}(X)\right|\leq c_{1}|\mathsf{p}(C)|^{\beta},X\in C).$
Summing over $C\in\mathcal{B}_{M}$ gives
$\sum_{C\in\mathcal{B}_{M}}\mathbb{E}[r_{T}^{\textrm{live}}(C)\mathbf{1}(\mathcal{G}_{C})]$ $\leq\sum_{C\in\mathcal{B}_{M}}c_{1}|\mathsf{p}(C)|^{\beta}(T-t_{M-1})\mathbb{P}(0<\left|f_{1}(X)-f^{(-1)}(X)\right|\leq c_{1}|\mathsf{p}(C)|^{\beta},X\in C)$
$=c_{1}w_{M-1}^{\beta}(T-t_{M-1})\mathbb{P}(0<\left|f_{1}(X)-f^{(-1)}(X)\right|\leq c_{1}w_{M-1}^{\beta})$ (46)
$\leq c_{1}w_{M-1}^{\beta}(T-t_{M-1})D_{0}\cdot\left[c_{1}w_{M-1}^{\beta}\right]^{\alpha}$
$=D_{0}(T-t_{M-1})[c_{1}w_{M-1}^{\beta}]^{1+\alpha}.$
Putting things together.
By combining the bounds of $V_{i},U_{i}$ and the regret of the last batch, we obtain
$R_{T}(\hat{\Gamma}_{\bm{u}},\hat{\pi}_{\bm{u}})$ $=\sum_{1\leq i<M}\left(U_{i}+V_{i}\right)+\sum_{C\in\mathcal{B}_{M}}\mathbb{E}[r_{T}^{\textrm{live}}(C)\mathbf{1}(\mathcal{G}_{C})]$
$\leq c\left(t_{1}+\sum_{i=2}^{M-1}(t_{i}-t_{i-1})\cdot w_{i-1}^{\beta+\alpha\beta}+(T-t_{M-1})w_{M-1}^{\beta+\alpha\beta}\right),$
where $c$ is a constant that depends on $(\beta,d)$.
Finally, we deal with the case of $\alpha=\infty$. By relations (44) and (D.2), one has
$U_{i}+V_{i}\leq(1+4\bar{c})c_{1}w_{i-1}^{\beta}(t_{i}-t_{i-1})\mathbb{P}(0<\left|f_{1}(X)-f^{(-1)}(X)\right|\leq c_{1}w_{i-1}^{\beta}).$ (47)
When $i=1$, the above relation simplifies to $U_{1}+V_{1}\leq(1+4\bar{c})c_{1}t_{1}$ because the probability term is upper-bounded by 1.
For $i\geq 2$, one has $c_{1}w_{i-1}^{\beta}<\delta_{0}$ due to the definition of $w_{i-1}$ and by the margin condition with $\alpha=\infty$, we get $\mathbb{P}(0<\left|f_{1}(X)-f^{(-1)}(X)\right|\leq c_{1}w_{i-1}^{\beta})=0$. Consequently, relation (47) reduces to $U_{i}+V_{i}=0$.
For the last batch, relation (D.2) similarly reduces to 0. Hence,
$R_{T}(\hat{\Gamma}_{\bm{u}},\hat{\pi}_{\bm{u}})\leq(1+4\bar{c})c_{1}t_{1}.$